### v2

# Random coverage from within with variable radii, and Johnson-Mehl cover timesthanks: Supported by EPSRC grant EP/T028653/1

Abstract
Given a compact planar region $A$, let $\tau_{A}$ be the (random) time it takes for $A$ to be fully covered by a spatial birth-growth process in $A$ with seeds arriving as a unit-intensity Poisson point process in $A\times[0,\infty)$, where upon arrival each seed grows at unit rate in all directions. We show that if $\partial A$ is smooth or polygonal then $\mathbb{P}[\pi\tau_{sA}^{3}-6\log s-4\log\log s\leq x]$ tends to $\exp(-(\frac{81}{4\pi})^{1/3}|A|e^{-x/3}\linebreak-(\frac{9}{2\pi^{2}})^{1/3}|\partial A|e^{-x/6})$ in the large-$s$ limit; the second term in the exponent is due to boundary effects, the importance of which was not recognized in earlier work on this model. We present similar results in higher dimensions (where boundary effects dominate). These results are derived using new results on the asymptotic probability of covering $A$ with a high-intensity spherical Poisson Boolean model restricted to $A$ with grains having iid small random radii, which generalize recent work of the first author that dealt only with grains of deterministic radius.
1 Introduction
The Johnson-Mehl (J-M) tessellation is a classic model of a random tessellation in $\mathbb{R}^{d}$, where $d\in\mathbb{N}$. Seeds are generated as a homogeneous Poisson point process ${\cal H}_{\rho}=\{(x_{i},t_{i})\}_{i\geq 1}$ of intensity $\rho$ in space-time $\mathbb{R}^{d}\times[0,\infty)$. If location $x_{i}$ is not already claimed by time $t_{i}$, the seed $i$ becomes a cell at that instant, which immediately starts to grow from $x_{i}$ at unit rate in all directions, claiming previously unclaimed territory as part of that cell. Whenever the growing cell hits another cell, it stops growing in that direction. Ultimately the whole of $\mathbb{R}^{d}$ is tessellated by cells.
This model dates back at least to work of Kolmogorov and of Johnson and Mehl in the 1930s on modelling crytallization processes; other applications include growth of surface film on metals, and more recent applications in neurobiology are described in [2]. For further discussion and references, see also [7, 8].
A natural variant is the restricted Johnson-Mehl tessellation, which we define as follows. Given a specified compact region $A\subset\mathbb{R}^{d}$, we allow only for seeds that arrive inside the region $A$, that is, the seeds are generated by a Poisson process in $A\times[0,\infty)$. Cells grow by the same rules as described before, and the restrictions of the cells to $A$ ultimately tessellate $A$.
We are interested here in the time at which $A$ becomes completely covered, i.e. the first time at which every point of $A$ has been claimed by a cell, either for the J-M model or for the restricted J-M model. In particular we are concerned with the limiting distribution of the cover time, denoted $\tilde{T}_{\rho}$ for the J-M model and $T_{\rho}$ for the restricted J-M model, as $\rho\to\infty$. Equivalently, one can keep $\rho$ fixed (say, $\rho=1$), and consider the distribution of the cover times for an expanding sequence of windows $(A_{L})_{L>0}$ given by dilations of a fixed set $A$ by a factor of $L$, in the large-$L$ limit. In the later case, the restricted Johnson-Mehl tessellation uses only seeds that arrive inside $A_{L}$, and we refer to the cover times in this limiting regime as $\tau_{L}$ (for the restricted J-M model) and $\tilde{\tau}_{L}$ (for the J-M model).
These cover times have previously been considered by S. N. Chiu in [1], for $A=[0,1]^{d}$. In fact, Chiu considers a more general class of J-M models, where the rate at which seeds arrive is allowed to be non-homogeneous in time, and where moreover the seeds do not necessarily all grow at the same rate. We do not consider such generalizations here.
In [1, page 893], Chiu rightly distinguishes between the J-M and restricted J-M models. However, he goes on to assert that ‘all theorems in this paper are valid for both models’ because ‘almost surely edge effects do not play a role in the limiting behaviour of $T_{L}$’ (Chiu’s $T_{L}$ is equivalent to our $\tau_{L}$). In other words, Chiu seems to be asserting that $\tau_{L}$ and $\tilde{\tau}_{L}$ have the same limiting distribution.
It is our contention that this assertion is incorrect. In dimensions $d\geq 2$, the limiting behaviour of $T_{\rho}$ is different from that of $\tilde{T}_{\rho}$, and the limiting behaviour of $\tau_{L}$ is different from that of $\tilde{\tau}_{L}$; in other words, edge effects do play a role. We justify this assertion with results identifying the limiting distribution of $T_{\rho}$ (suitably scaled and centred) and showing it is different from that of $\tilde{T}_{\rho}$; our limiting result for $\tilde{T}_{\rho}$ is consistent with the result in [1]. Likewise we show that the limiting distribution of $\tau_{L}$ (suitably scaled and centred) is different from that of $\tilde{\tau}_{L}$. Our results for $T_{\rho}$ and $\tau_{L}$ apply when $A$ is polygonal in $d=2$, or when it has a smooth (more precisely, $C^{2}$) boundary for general $d\geq 2$.
When $A=[0,1]^{d}$, we do not provide detailed limiting distributions for $T_{\rho}$ and $\tau_{L}$ except in the case $d=2$. To give a detailed limit distribution when $d\geq 3$ would require careful consideration of all faces of all dimensions and is beyond the scope of this paper. However, the time to cover all the $(d-1)$-dimensional faces will be a lower bound for the actual cover time so boundary effects can certainly not be neglected in this case either.
We shall derive our results on the J-M model from new results, of independent interest, on the spherical Poisson Boolean model (SPBM), which is defined to be a collection of Euclidean balls (referred to as grains) of i.i.d. random radii, centred on the points of a homogeneous Poisson process in the whole of $\mathbb{R}^{d}$. For the the restricted SPBM, we take a Poisson process on $A$ rather than on all of $\mathbb{R}^{d}$.
We shall determine the probability that the restricted SPBM covers the whole of $A$, in the limit when the Poisson intensity becomes large and the radii of balls become small in a linked manner. In [10] we derived results of this nature for balls with deterministic radius; here we generalize them to allow for balls of random radius, which is needed to derive our results on the J-M cover time. Our results on coverage of $A$ by the restricted SPBM complement the classic results of [3, 5] on the limiting probability of covering $A$ with an unrestricted SPBM.
2 Statement of results
Throughout this paper, we assume that we are given $d\in\mathbb{N}$, and a compact, Riemann measurable set $A\subset\mathbb{R}^{d}$ (Riemann measurability of a bounded set in $\mathbb{R}^{d}$ amounts to its boundary having zero Lebesgue measure).
[…]
Define the set $D^{[r]}$ to be the interior of the union of all hypercubes of the form $\prod_{i=1}^{d}[n_{i}r,(n_{i}+1)r]$, with $n_{1},\ldots,n_{d}\in\mathbb{Z}$, that are contained in $D^{o}$ (the set $D^{[r]}$ resembles $D^{(r)}$ but is guaranteed to be Riemann measurable).
We say that $D$ has $C^{2}$ boundary (for short: $\partial D\in C^{2}$) if for each $x\in\partial D$ there exists a neighbourhood $U$ of $x$ and a real-valued function $f$ that is defined on an open set in $\mathbb{R}^{d-1}$ and twice continuously differentiable, such that $\partial D\cap U$, after a rotation, is the graph of the function $f$. We say that $\partial D\in C^{1,1}$ (a weaker condition) if for each $x$ the function $f$ satisfies only that $f$ is contiuously differentiable with Lipschitz first order partial derivatives.
2.1 Results for the spherical Poisson Boolean model
As discsussed in the introduction, we shall derive our results for the J-M model from results on coverage by the restricted SPBM, which are of independent interest and which we now state. Given a nonnegative random variable $Y$ and given $n>0$, suppose we have a collection of balls of independent random radii with the distribution of $Y$, centred on the points of a homogeneous Poisson point process in $A$ of intensity $n$.
Given also $k\in\mathbb{N}$, let $Z_{A,k}(n,Y)$ be the set of points $x\in\mathbb{R}^{d}$ such that $x$ is covered by at least $k$ of the balls, so in particular $Z_{A,1}(n,Y)$ is the union of the balls (we shall provide a more detailed description of $Z_{A,k}(n,Y)$ in Section 3.2.) Throughout this paper, $n$ is not necessarily an integer.
Theorem 2.1.
Suppose $Y$ is a nonnegative random variable with $0<\mathbb{E}\,[Y^{2+\varepsilon}]<\infty$ for some $\varepsilon>0$. Suppose that $d=2$ and $A$ is polygonal. Let $k\in\mathbb{N},\beta\in\mathbb{R}$, and suppose $(r_{n})_{n>0}$ are nonnegative numbers satisfying $n\pi r_{n}^{2}\mathbb{E}\,[Y^{2}]-\log n-(2k-1)\log\log n\to\beta$ as $n\to\infty$. Then as $n\to\infty$,
$\mathbb{P}[A\subset Z_{A,k}(n,r_{n}Y)]\to\exp\Big{(}-\Big{(}\frac{(\mathbb{E}\,[Y])^{2}}{\mathbb{E}\,[Y^{2}]}\Big{)}{\bf 1}_{\{k=1\}}|A|e^{-\beta}-\Big{(}\frac{c_{2,k}\mathbb{E}\,[Y]|\partial A|}{(\mathbb{E}\,[Y^{2}])^{1/2}}\Big{)}e^{-\beta/2}\Big{)},$ (2.1)
where we set $c_{2,k}:=\pi^{-1/2}(1/2)^{k-1}/(k-1)!$.
Given $d\in\mathbb{N}$, let $\omega_{d}:=\pi^{d/2}/\Gamma(1+d/2)$, the volume of the unit ball in $\mathbb{R}^{d}$, and set
$c_{d}:=\frac{1}{d!}\left(\frac{\sqrt{\pi}\;\Gamma(1+d/2)}{\Gamma((d+1)/2)}\right)^{d-1}.$ (2.2)
Note that $c_{1}=c_{2}=1$, and $c_{3}=3\pi^{2}/32$. Moreover, using Stirling’s formula one can show that $c_{d}^{1/d}\sim e(\pi/(2d))^{1/2}$ as $d\to\infty$. The constant $c_{d}$ is denoted $\psi_{d}$ in [1, page 894], while $\omega_{d}$ is denoted $\theta_{d}$ in [10]. Later it will be useful to have $\omega_{0}$ defined: we set $\omega_{0}:=1$.
Our result for $A$ with $\partial A\in C^{2}$ involves a further constant $c_{d,k}$ defined (as in [10]) for each $d,k\in\mathbb{N}$ with $d\geq 2$ by
$c_{d,k}$ $:=c_{d-1}\omega_{d}^{2-d-1/d}\omega_{d-1}^{2d-3}\omega_{d-2}^{1-d}(1-1/d)^{d+k-3+1/d}2^{-1+1/d}/(k-1)!.$ (2.3)
It can be checked that
$c_{d,1}=((d-1)!)^{-1}2^{1-d}(d-1)^{d-2+1/d}\pi^{(d/2)-1}\Gamma\left(\frac{d+1}{2}\right)^{1-d}\Gamma\left(\frac{d}{2}\right)^{d-1+1/d},$ (2.4)
and $c_{d,k}=c_{d,1}(1-1/d)^{k-1}/(k-1)!$. By (2.3) we have $c_{2,1}=\pi^{-1/2}$ and $c_{3,1}=2^{-4}\pi^{5/3}$. In the right hand side of (2.4) the first factor of $((d-1)!)^{-1}$ was given, incorrectly, as $(d!)^{-1}$ in [10].
Theorem 2.2.
Suppose that $d\geq 2$ and $\partial A\in C^{2}$ with $\overline{A^{o}}=A$. Suppose $Y$ is a nonnegative random variable with $0<\mathbb{E}\,[Y^{d+\varepsilon}]<\infty$ for some $\varepsilon>0$. Let $k\in\mathbb{N},\beta\in\mathbb{R}$, and suppose $(r_{n})_{n>0}$ are nonnegative numbers satisfying
$n\omega_{d}r_{n}^{d}\mathbb{E}\,[Y^{d}]-(2-2/d)\log n-2(d+k-3+1/d)\log\log n\to\beta$ (2.5)
as $n\to\infty$. If $d=2$ then (2.1) holds, while if $d\geq 3$ then
$\lim_{n\to\infty}\mathbb{P}[A\subset Z_{A,k}(n,r_{n}Y)]=\exp\Big{(}-c_{d,k}\Big{(}\frac{(\mathbb{E}\,[Y^{d-1}])^{d-1}}{(\mathbb{E}\,[Y^{d}])^{d-2+1/d}}\Big{)}|\partial A|e^{-\beta/2}\Big{)}.$ (2.6)
Remark 2.3.
Taking $Y\equiv 1$ in these results gives [10, Theorem 3.2] and [10, Theorem 3.1]. [10] also includes parallel results (in the case $Y\equiv 1$) for a situation where the number of grains in the spherical Boolean model is a large deterministic constant rather than a Poisson variable. We expect that a result along these lines could be given for general $Y$ using the methods of this paper.
Remark 2.4.
The condition $\overline{A^{o}}=A$ should have been included in the statement of [10, Theorem 3.1]. For example, if $d=2$ and $A$ is the union of a disk (filled in) and a circle (not filled in, and disjoint from the disk) then that theorem does not apply.
Remark 2.5.
In Theorem 2.2, we conjecture that the condition $\partial A\in C^{2}$ can be relaxed to $\partial A\in C^{1,1}$.
Remark 2.6.
If $\mathbb{E}\,[Y^{d}]=\infty$ then $\mathbb{P}[A\subset Z_{\mathbb{R}^{d},k}(n,rY)]=1$ for any $k\in\mathbb{N}$ and any $r>0$; see e.g. [6, Theorem 16.4] for the case $k=1$. However, one could still look for a sequence $(r_{n})$ such that $\mathbb{P}[A\subset Z_{A,1}(n,r_{n}Y)]$ converges to a nontrivial limit. It seems plausible in this case that $A$ is most likely to be covered (if at all) by a small number of balls. This looks like an interesting problem that lies beyond the scope of this paper.
Remark 2.7.
A series of videos illustrating the SPBM and Remark 2.6 can be viewed at this link [^2] https://www.youtube.com/playlist?list=PLiaV5rk6Gk7qEXpLOU7FSvN4b8dy1_GJn. There are three videos, with $n=1000$, $n=10000$ and $n=20000$ respectively. Each video has a Poisson point process of intensity $n$ inside $[0,1]^{2}$ with points $\{p_{1},\dots,p_{N_{n}}\}$. The frame at time $t$ is an image of a tessellation corresponding to a restricted SPBM, $Z_{A,1}(n,r_{t}Y_{t})$, where $A=[0,1]^{2}$, $Y_{t}$ is Pareto distributed with shape parameter $\alpha(t)$ and $r_{t}$ is chosen so that $A\subset Z_{A,1}(n,r_{t}Y_{t})$. There are $N_{n}$ cells in the frame: the pixel with centre $x\in[0,1]^{2}$ is assigned to a cell based on $\mathrm{argmin}_{i}\inf\{r\geq 0:x\in B(p_{i},rY_{t,i})\}$, where $(Y_{t,i})_{1\leq i\leq N_{n}}$ are independent copies of $Y_{t}$. In other words, for $t$ fixed the tessellation is generated by a growth process for which all seeds are born at time 0, growth rates are random and distributed like $Y_{t}$, and the growing cells are interpenetrable.
Note that cells are not necessarily connected. When displaying the video we colour the cells so that no two adjacent cells have the same colour, although two distinct cells which never touch are allowed to have the same colour.
Over the video, $\alpha(t)$ changes from $3.00$ to $0.25$, i.e. $Y_{t}$ becomes more heavy-tailed, with the $Y_{t}$s coupled using the definition $Y_{t}:=U^{-1/\alpha(t)}$ where $U$ is uniformly distributed on $[0,1]$.
At the beginning of each video, $\mathbb{E}\,[Y_{t}^{3-\varepsilon}]<\infty$ for all $\varepsilon>0$ so Theorem 2.1 applies. As the video progresses, we gradually reduces the number of finite moments, finishing when $\mathbb{E}\,[Y_{t}^{0.25}]=\infty$.
The videos after the point where $\mathbb{E}\,[Y_{t}^{2}]=\infty$ illustrate Remark 2.6. In the period when $\mathbb{E}\,[Y_{t}^{2}]=\infty$ but $\mathbb{E}\,[Y_{t}^{7/4}]<\infty$, there is not yet a single cell containing more than half the area of $[0,1]^{2}$, indicating that there may be a non-trivial period where several balls are required to cover $[0,1]^{2}$.
2.2 Results for the Johnson-Mehl model
Assume that on a common probability space $(\mathbb{S},{\cal F},\mathbb{P})$, we have a family of Poisson point processes ${\cal H}_{\rho},\rho>0$. Here ${\cal H}_{\rho}$ is homogeneous of intensity $\rho$ on $\mathbb{R}^{d}\times\mathbb{R}_{+}$, where $\mathbb{R}_{+}:=[0,\infty)$. We write ${\cal H}_{\rho,A}$ for ${\cal H}_{\rho}\cap(A\times\mathbb{R}_{+})$.
Given $\rho\in(0,\infty)$, define the coverage times $T_{\rho}$ and $\tilde{T}_{\rho}$ by
$T_{\rho}:=\inf\left\{t>0:A\subset\cup_{(x,s)\in{\cal H}_{\rho}\cap(A\times[0,t])}B(x,t-s)\right\};$ (2.7)
$\tilde{T}_{\rho}:=\inf\left\{t>0:A\subset\cup_{(x,s)\in{\cal H}_{\rho}\cap(\mathbb{R}^{d}\times[0,t])}B(x,t-s)\right\}.$ (2.8)
Also, for $L>0$ set $A_{L}:=\{Lx:x\in A\}$ and let
$\tau_{L}:=\inf\left\{t>0:A_{L}\subset\cup_{(x,s)\in{\cal H}_{1}\cap(A_{L}\times[0,t])}B(x,t-s)\right\};$ (2.9)
$\tilde{\tau}_{L}:=\inf\left\{t>0:A_{L}\subset\cup_{(x,s)\in{\cal H}_{1}\cap(\mathbb{R}^{d}\times[0,t])}B(x,t-s)\right\}.$ (2.10)
Then $T_{\rho}$, $\tilde{T}_{\rho}$ are the cover times of $A$ for the restricted J-M model and for the J-M model, respectively, with intensity $\rho$. Likewise $\tau_{L}$, $\tilde{\tau}_{L}$, are the cover time of $A_{L}$ for the restricted J-M model and for the J-M model, respectively, with intensity $1$.
Our results in this subsection are concerned with the asymptotic distribution of the cover times $T_{\rho}$ and $\tilde{T}_{\rho}$ as $\rho\to\infty$, and the asymptotic distribution of $\tau_{L}$ and $\tilde{\tau}_{L}$ as $L\to\infty$. We start with a result in general $d$ for the unrestricted cover time $\tilde{T}_{\rho}$ (and for $\tilde{\tau}_{L})$. This is simpler to deal with than $T_{\rho}$ because boundary effects are avoided.
Proposition 2.8.
Suppose $A\subset\mathbb{R}^{d}$ is compact and Riemann measurable with $|A|>0$. Let $\beta\in\mathbb{R}$. Then as $\rho\to\infty$,
$\mathbb{P}\left[\omega_{d}\rho\tilde{T}_{\rho}^{d+1}-d\log\rho-d^{2}\log\log\rho\leq\beta\right]\to\exp(-c_{d}(d^{d}\omega_{d})^{-1/(d+1)}|A|e^{-\beta/(d+1)}).$ (2.11)
Also as $L\to\infty$,
$\mathbb{P}\left[\omega_{d}\tilde{\tau}_{L}^{d+1}-d(d+1)\log L-d^{2}\log\log L\leq\beta\right]$
$\to\exp\Big{(}-c_{d}\Big{(}\frac{(d+1)^{d^{2}}}{d^{d}\omega_{d}}\Big{)}^{1/(d+1)}|A|e^{-\beta/(d+1)}\Big{)}.$ (2.12)
In Appendix A, we shall verify that this result is consistent with the case $\gamma=v=1$ of [1, Theorem 4]. The proof in [1] refers to an unpublished preprint; for completeness we shall present the (fairly short) proof of Proposition 2.8 in Section 4.
Our main new results in this section are concerned with $T_{\rho}$ and $\tau_{L}$, the cover times for the restricted J-M model. We first give a result for the case where $A$ is a polygon in $d=2$.
Theorem 2.9.
Suppose that $d=2$ and $A$ is polygonal. Let $\beta\in\mathbb{R}$. Then as $\rho\to\infty$,
$\mathbb{P}[\pi\rho T_{\rho}^{3}-2\log\rho-4\log\log\rho\leq\beta]\to\exp(-(4\pi)^{-1/3}|A|e^{-\beta/3}-(2\pi^{2})^{-1/3}|\partial A|e^{-\beta/6}).$ (2.13)
Also as $L\to\infty$,
$\mathbb{P}[\pi\tau_{L}^{3}-6\log L-4\log\log L\leq\beta]\to\exp\Big{(}-\Big{(}\frac{81}{4\pi}\Big{)}^{1/3}|A|e^{-\beta/3}-\Big{(}\frac{9}{2\pi^{2}}\Big{)}^{1/3}|\partial A|e^{-\beta/6}\Big{)}.$ (2.14)
Our result for general $d$ and $A$ with $C^{2}$ boundary involves the constant $c^{\prime}_{d}$ defined by
$c^{\prime}_{d}$ $:=\frac{c_{d-1}\omega_{d-1}^{2d-3}}{\omega_{d-2}^{d-1}d^{d-1}}\Big{(}\frac{(d-1)^{d}}{2\omega_{d}^{d}}\Big{)}^{(d-1)/(d+1)}.$ (2.15)
Theorem 2.10.
Suppose that $d\geq 2$ and $\partial A\in C^{2}$ with $\overline{A^{o}}=A$. Let $\beta\in\mathbb{R}$. Then as $\rho\to\infty$, if $d=2$ then (2.13) holds while if $d\geq 3$ then
$\mathbb{P}[\omega_{d}\rho T_{\rho}^{d+1}-2(d-1)\log\rho-2d(d-1)\log\log\rho\leq\beta]\to\exp(-c^{\prime}_{d}|\partial A|e^{-\beta/(2d+2)}).$ (2.16)
Also as $L\to\infty$, if $d=2$ then (2.14) holds while if $d\geq 3$ then
$\mathbb{P}[\omega_{d}\tau_{L}^{d+1}-2(d^{2}-1)\log L-2d(d-1)\log\log L\leq\beta]$
$\to\exp(-(d+1)^{d(d-1)/(d+1)}c^{\prime}_{d}|\partial A|e^{-\beta/(2d+2)}).$ (2.17)
Figure 1: Empirical distributions of the standardised coverage times $T_{\rho}$ and $\tilde{T}_{\rho}$ as described in Proposition 2.8 and Theorem 2.9, obtained by sampling many independent realisations of the Johnson-Mehl process in $[0,1]^{2}$. The standardisation function $g(T,\rho)=\pi\rho T^{3}-2\log\rho-4\log\log\rho$ is the same for both $T_{\rho}$ and $\tilde{T}_{\rho}$. In [1], Chiu proves a weak limit theorem for $\tilde{T}_{\rho}$ (or rather, for $\tilde{\tau}_{L}$, which is equivalent to a limit theorem for $\tilde{T}_{\rho}$) with a different standardisation. In the large-$\rho$ limit, we show in Appendix A that our weak law in Proposition 2.8 is consistent with [1]. Chiu suggested that the same limit theorem should apply to $\tau_{L}$ as to $\tilde{\tau}_{L}$, as the boundary effects would not affect the limit. Our Theorem 2.9, and the above figure, show that in fact the boundary effects do affect the limiting distribution of $g(T_{\rho},\rho)$.
In short, when $d=2$, for $\rho$ large the last point of $A$ to get covered in the restricted J-M model could lie either on the boundary or in the interior of $A$, while for $d\geq 3$, when $\rho$ is large the last point to get covered is very likely to lie on the boundary of $A$.
Remark 2.11.
A series of videos of very high intensity restricted Johnson-Mehl processes inside $A=[0,1]^{2}$ can be viewed here [^3] https://www.youtube.com/playlist?list=PLiaV5rk6Gk7rh6nvAts1w4WZrHE0TKeW2 An interactive tool for generating Johnson-Mehl tessellations is available at https://frankiehiggs.pyscriptapps.com/johnson-mehl-plot/latest. The cells are coloured (with five different colours) so that no two adjacent cells in the final “tessellation” have the same colour.
To illustrate the above remark about which point is covered last when $d=2$, we have included one video in the playlist where the last point to be covered is at the boundary, and one where it is in the interior.
Results of more simulations, comparing the distribution of cover times for the restricted and unrestricted J-M models with finite $\rho$ and their limiting distributions, are presented in Figure 1.
Remark 2.12.
One could consider a slight variant of the restricted J-M model, in which a growing cell, whenever it hits the boundary of $A$, stops growing in that direction. Let $T^{*}_{\rho}$ denotes the time at which the union of cells first covers $A$ for this variant of the model.
When $A$ is convex, the value of $T^{*}_{\rho}$ is the same as that of $T_{\rho}$. If $A$ is not convex, we would expect our results to still be true for $T^{*}_{\rho}$ as well as for $T_{\rho}$, but we do not prove this.
Remark 2.13.
Suppose we take $d=2$ and $A=[0,1]^{2}$. Comparing (2.11) with (2.13) we see that the limiting distribution of $T_{\rho}^{3}$ (appropriately scaled and centred) is different from that of $\tilde{T}_{\rho}^{3}$ (scaled and centred the same way). Similarly, comparing (2.12) with (2.14) we see that the limiting distribution of $\tau_{L}^{3}$ (appropriately scaled and centred) is different from that of $\tilde{\tau}_{L}^{3}$ (scaled and centred the same way).
When $d\geq 3$ and $\partial A\in C^{2}$, we see from (2.11) and (2.16) that even the centring constants required to get a nondegenerate limit distribution for $\omega_{d}\rho T_{\rho}^{d+1}$ are different (larger) compared to the corresponding constants for $\omega_{d}\rho\tilde{T}_{\rho}^{d+1}$. Again similar remarks apply for $\tau_{L}$ and $\tilde{\tau}_{L}$.
Remark 2.14.
As mentioned earlier, [1] considers cover times for more general J-M models where the growth rate of cells, or the arrival rates of seeds, are not necessarily constants. If growth rates are non-constant, then it is rather problematic to model the covered region at a given time as simply a SPBM; one has to modify our initial description of the cell growth model to allow a faster-growing cell to pass through a slower-growing cell and come out the other side. If one is willing to do this (which seems to be the approach in [1]), the methods of this paper should be applicable for dealing with boundary effects for these more general J-M models too.
Remark 2.15.
The distribution of the maximum of two independent Gumbel variables with different scale parameters is known as a two-component extreme value (TCEV) distribution in the hydrology literature [12].
Theorems 2.9 and 2.10 tell us that when $d=2$ and $A$ is polygonal or $\partial A\in C^{2}$, the random variables $\omega_{d}\rho T_{\rho}^{d+1}$ and of $\omega_{d}\tau_{L}^{d+1}$, suitably centred, are asymptotically TCEV distributed, while if $d\geq 3$ and $\partial A\in C^{2}$, they are asymptotically Gumbel distributed.
2.3 Strategy of proof
Next, we briefly describe the strategy for the proof, in Sections 3 and 4, of the weak convergence results that were stated above.
We shall use a known result (Lemma 3.2) giving the limiting probability of covering a bounded region of $\mathbb{R}^{d}$ by an SPBM on the whole of $\mathbb{R}^{d}$, in the limit of high intensity $n$ and small balls of random radius, i.e. distributed as $r_{n}Y$ for a specified nonnegative random variable $Y$ and for constants $r_{n}$ that become small as $n$ becomes large.
We shall consider for large $n$ and suitable choice of $r_{n}$, the SPBM with balls of radii distributed as $r_{n}Y$, restricted to a $d$-dimensional half-space $\mathbb{H}$. In Lemma 3.5 we determine the limiting probability that a given bounded set within the hyperplane $\partial\mathbb{H}$ is covered, by applying Lemma 3.2 in $d-1$ dimensions. Moreover we will show that the probability that a region in the half-space within distance $n^{\zeta}r_{n}$ of that set is covered with the same limiting probability, where ${\zeta}$ is a small positive constant.
We shall then prove Theorem 2.1 by applying Lemma 3.5 to determine the limiting probability that the region near the edges of a polygonal set $A$ is covered, and Lemma 3.2 directly to determine the limiting probability that the interior region is covered, along with a separate argument to show the regions near the corners of $A$ are covered with high probability.
To prove Theorem 2.2 we approximate to $A$ by a polytopal set $A_{n}$ with faces of width $O(n^{9{\zeta}}r_{n})$ and follow similar steps to those just mentioned for polygonal $A$.
In the case where $Y$ is unbounded, to obtain the required independence, for example between coverage events for different faces of our polyhedron or polygon, we consider only those balls of radius at most $n^{\zeta}r_{n}$, with a separate argument (Lemmma 3.3) to show this does not affect the limiting coverage probabilities.
Applying Lemma 3.2 for an SPBM with $Y$ uniformly distributed on $[0,1]$ yields a proof of Proposition 2.8. Similarly, applying Theorems 2.1 and 2.2 for an SPBM with $Y$ uniformly distributed on $[0,1]$ yields proofs of Theorems 2.9 and 2.10 respectively.
3 Proof of Theorems 2.1 and 2.2
Throughout this section, we assume $d,k\in\mathbb{N}$ are fixed with $d\geq 2$, and $A\subset\mathbb{R}^{d}$ is compact and nonempty with $A=\overline{A^{o}}$, and that we are given a nonnegative random variable $Y$ satisfying $0<\mathbb{E}\,[Y^{\gamma}]<\infty$ for some $\gamma>d$.
3.1 Preliminaries
We now give some further notation used throughout. Let $o$ denote the origin in $\mathbb{R}^{d}$. Set $\mathbb{H}:=\mathbb{R}^{d-1}\times[0,\infty)$ and $\partial\mathbb{H}:=\mathbb{R}^{d-1}\times\{0\}$.
Given two sets $\mathcal{X},{\cal Y}\subset\mathbb{R}^{d}$, we set $\mathcal{X}\triangle{\cal Y}:=(\mathcal{X}\setminus{\cal Y})\cup({\cal Y}\setminus\mathcal{X})$, the symmetric difference between $\mathcal{X}$ and ${\cal Y}$. Also, we write $\mathcal{X}\oplus{\cal Y}$ for the set $\{x+y:x\in\mathcal{X},y\in{\cal Y}\}$, and #$\#(\mathcal{X})$ for the number of elements of $\mathcal{X}$ (possibly +$\infty$+ ∞). Given also $x\in\mathbb{R}^{d}$ we write $x+{\cal Y}$ for $\{x\}\oplus{\cal Y}$.
Given a Borel measure $\mu$ on $\mathbb{R}^{d}$, and Borel $D\subset\mathbb{R}^{d}$, let $\mu|_{D}$ denote the restriction of the measure $\mu$ to $D$ (a Borel measure on $D$).
Given $x,y\in\mathbb{R}^{d}$, we denote by $[x,y]$ the line segment from $x$ to $y$, that is, the convex hull of the set $\{x,y\}$. We write $a\wedge b$ (respectively $a\vee b$) for the minimum (resp. maximum) of any two numbers $a,b\in\mathbb{R}$.
Given $m\in\mathbb{N}$ and functions $f:\mathbb{N}\cap[m,\infty)\to\mathbb{R}$ and $g:\mathbb{N}\cap[m,\infty)\to(0,\infty)$, we write $f(n)=O(g(n))$ as $n\to\infty$ if $\limsup_{n\to\infty}|f(n)|/g(n)<\infty$. We write $f(n)=o(g(n))$ as $n\to\infty$ if $\lim_{n\to\infty}f(n)/g(n)=0$, and $f(n)\sim g(n)$ as $n\to\infty$ if $\lim_{n\to\infty}f(n)/g(n)=1$. We write $f(n)=\Theta(g(n))$ as $n\to\infty$ if $f$ takes only positive values, and both $f(n)=O(g(n))$ and $g(n)=O(f(n))$. Given $s>0$ and functions $f:(0,s)\to\mathbb{R}$ and $g:(0,s)\to(0,\infty)$, we write $f(r)=O(g(r))$ as $r\downarrow 0$ $\limsup_{r\downarrow 0}|f(r)|/g(r)<\infty$, and write $g(r)=\Omega(g(r))$ as $r\downarrow 0$, if $\liminf_{r\downarrow 0}f(r)/g(r)>0$. We write $f(r)=o(g(r))$ as $r\downarrow 0$ if $\lim_{r\downarrow 0}f(r)/g(r)=0$, and $f(r)\sim g(r)$ as $r\downarrow 0$ if this limit is 1.
From time to time we shall use the following geometrical lemma.
Lemma 3.1 (Geometrical lemma).
Suppose $\partial A\in C^{2}$, and $A=\overline{A^{o}}$. Given $\varepsilon>0$, there exists $r_{0}=r_{0}(d,A,\varepsilon)>0$ such that
$|B(x,r)\cap A|\geq((\omega_{d}/2)-\varepsilon)r^{d},~{}~{}~{}~{}\forall x\in A,r\in(0,r_{0}).$ (3.1)
Proof.
See [4, Lemma 3.2(a)]
We shall repeatedly use the following result from [10], which is based on results in Janson [5] or Hall [3]. Recall that $c_{d}$ and $Z_{A}(n,Y)$ were defined at (2.2) and just before Theorem 2.1 respectively.
Lemma 3.2 (Limiting probability of coverage by an unrestricted SPBM).
Let $d\in\mathbb{N}$. Let $\alpha:=\omega_{d}\mathbb{E}\,[Y^{d}]$. Let $\beta\in\mathbb{R}$. Suppose $\delta(\lambda)\in(0,\infty)$ is defined for all $\lambda>0$, and satisfies
$\lim_{\lambda\to\infty}\left(\alpha\delta(\lambda)^{d}\lambda-\log\lambda-(d+k-2)\log\log\lambda\right)=\beta.$ (3.2)
Let $B\subset\mathbb{R}^{d}$ be compact and Riemann measurable, and for each $\lambda>0$ let $B_{\lambda}\subset B$ be Riemann measurable with the properties that $B_{\lambda}\subset B_{\lambda^{\prime}}$ whenever $\lambda\leq\lambda^{\prime}$, and that $\cup_{\lambda>0}B_{\lambda}\supset B^{o}$. Then
$\lim_{\lambda\to\infty}\mathbb{P}[B_{\lambda}\subset Z_{\mathbb{R}^{d},k}(\lambda,\delta(\lambda)Y)=\exp\left(-\left(\frac{c_{d}(\mathbb{E}\,[Y^{d-1}])^{d}}{(k-1)!(\mathbb{E}\,[Y^{d}])^{d-1}}\right)|B|e^{-\beta}\right).$ (3.3)
Proof.
See [10, Lemma 7.2]. There, it was assumed that $Y$ is bounded, but the proof carries over if this condition is relaxed to the $(d+\varepsilon)$-th moment condition used here. ∎
3.2 Coverage in the boundary by a SPBM in a half-space
In this subsection, we assume $(r_{n})_{n>0}$ satisfies (2.5) for some $\beta\in\mathbb{R}$. Then as $n\to\infty$,
$r_{n}=\left(\frac{(2-2/d)\log n+2(d+k-3+1/d)\log\log n+\beta+o(1)}{n\omega_{d}\mathbb{E}\,[Y^{d}]}\right)^{1/d},$ (3.4)
and
$\exp(-\omega_{d}nr_{n}^{d}\mathbb{E}\,[Y^{d}])\sim n^{-(2-2/d)}(\log n)^{-2(d+k-3+1/d)}e^{-\beta}.$ (3.5)
We shall use the following notation throughout the sequel. Given $n>0$ let ${\cal U}_{n}$ be a Poisson point process in $\mathbb{R}^{d}\times\mathbb{R}_{+}$ (viewed as a random set of points in $\mathbb{R}^{d+1}$) with intensity measure $n{\rm Leb}_{d}\otimes\mu_{Y}$, where ${\rm Leb}_{d}$ denotes $d$-dimensional Lebesgue measure and $\mu_{Y}$ denotes the distribution of $Y$. Given Borel $D\subset\mathbb{R}^{d}$, and given any point process $\mathcal{X}$ in $\mathbb{R}^{d}\times\mathbb{R}_{+}$, we define
${\cal U}_{n,D}$ $:={\cal U}_{n}\cap(D\times\mathbb{R}_{+});$ (3.6)
$Z_{n}(\mathcal{X})$ :=$:=\{y\in\mathbb{R}^{d}:\#\{(x,s)\in\mathcal{X}:y\in B(x,r_{n}s)\}\geq k\};$ (3.7)
$Z_{n}^{o}(\mathcal{X})$ $:=(Z_{n}(\mathcal{X}))^{o}.$ (3.8)
Thus, ${\cal U}_{n,D}$ is a Poisson process in $D\times\mathbb{R}_{+}$ with intensity measure $n{\rm Leb}_{d}\otimes\mu_{Y}$ (strictly speaking, with intensity measure $n{\rm Leb}_{d}|_{D}\otimes\mu_{Y}$). Equivalently, it can be viewed as a homogeneous, independently marked Poisson process in $D$ of intensity $n$ with marks having the distribution of $Y$ (see e.g. [6, Theorem 5.6].)
The set $Z_{n}({\cal U}_{n,D})$ is the region that is covered at least $k$ times (for short: the $k$-covered region) for the restricted SPBM in $D$ with intensity $n$ and with radii having the distribution of $r_{n}Y$. It is the same as $Z_{D,k}(n,r_{n}Y)$ in earlier notation; we now consider $k$ to be fixed and suppress it from the notation. The set $Z_{n}^{o}({\cal U}_{n,D})$ is the region covered at least $k$ times by a union of open balls.
Recall that $\mathbb{H}:=\mathbb{R}^{d-1}\times[0,\infty)$. The main result of this section is Lemma 3.5 below, which determines the limiting probability of covering a $(d-1)$-dimensional region of the form $\Omega\times\{0\}$ in the hyperplane $\partial\mathbb{H}:=\mathbb{R}^{d-1}\times\{0\}$ by $Z_{n}({\cal U}_{n,\mathbb{H}})$, or of covering the $a_{n}r_{n}$-neighbourhood of $\Omega\times\{0\}$ in $\mathbb{H}$, for $a_{n}$ not growing too fast with $n$. It is crucial for dealing with boundary regions in the proof of Theorem 2.10.
Before getting to that, we present another idea which will be important later. Let ${\zeta}>0$ to be chosen later (think of ${\zeta}$ as a small constant). Given Borel $D\subset\mathbb{R}^{d}$, we consider two separate, independent Poisson processes ${\cal U}^{\prime}_{n,D}$ and ${\cal U}^{\prime}_{n,D}$, defined by
${\cal U}^{\prime}_{n,D}={\cal U}_{n}\cap(D\times[0,n^{\zeta}]);~{}~{}~{}~{}~{}~{}~{}~{}~{}{\cal U}^{\prime\prime}_{n,D}={\cal U}_{n}\cap(D\times(n^{\zeta},\infty)).$ (3.9)
For various sets $D$ and ultimately for $D=A$, we shall be interested in coverage of certain regions by the set $Z_{n}({\cal U}_{n,D})$; the next lemma says that the limiting probability of coverage is unaffected if we take $Z_{n}({\cal U}^{\prime}_{n,D})$ instead of $Z_{n}({\cal U}_{n,D})$; this is useful because the random sets $Z_{n}({\cal U}^{\prime}_{n,D})$ have better spatial independence properties, considered as a function of $D$.
Lemma 3.3 (Asymptotic equivalence of coverage events using ${\cal U}_{n,D}$ or ${\cal U}^{\prime}_{n,D}$).
Let $n_{0}\in(0,\infty)$. Suppose that Borel $D_{n}\subset\mathbb{R}^{d}$ and $E_{n}\subset\mathbb{R}^{d}$ are defined for each $n\geq n_{0}$. Then as $n\to\infty$,
[$\mathbb{P}[\{E_{n}\subset Z_{n}({\cal U}_{n,D_{n}})\}\setminus\{E_{n}\subset Z_{n}({\cal U}^{\prime}_{n,D_{n}})\}]\to 0.$ (3.10)
Proof.
Recall that we are assuming $\mathbb{E}\,[Y^{\gamma}]<\infty$ for some $\gamma>d$. Given $x\in\mathbb{R}^{d}$, let $N_{n}(x)$ denote the number of points $(y,t)\in{\cal U}^{\prime\prime}_{n,D_{n}}$ such that $\|y-x\|\leq r_{n}t$. Then by the Hölder and Markov inequalities,
$\mathbb{P}[N_{n}(x)\geq 1]\leq\mathbb{E}\,[N_{n}(x)]$ $\leq n\int_{(n^{\zeta},\infty)}|B(x,r_{n}t)|\mu_{Y}(dt)$
$=n\omega_{d}r_{n}^{d}\mathbb{E}\,[Y^{d}{\bf 1}_{\{Y>n^{\zeta}\}}]$
$\leq n\omega_{d}r_{n}^{d}(\mathbb{E}\,[Y^{\gamma}])^{d/\gamma}(\mathbb{P}[Y>n^{\zeta}])^{(\gamma-d)/\gamma}$
$\leq\omega_{d}(\mathbb{E}\,[Y^{\gamma}])^{d/\gamma}nr_{n}^{d}(\mathbb{E}\,[Y^{\gamma}]/n^{{\zeta}\gamma})^{(\gamma-d)/\gamma},$
which tends to zero since $nr_{n}^{d}=O(\log n)$ by (3.4). Therefore
$\sup_{x\in\mathbb{R}^{d}}\mathbb{P}[x\in Z_{n}({\cal U}^{\prime\prime}_{n,D_{n}})]\to 0~{}~{}{\rm as}~{}n\to\infty.$ (3.11)
Let $n\geq n_{0}$. Since $E_{n}\subset\mathbb{R}^{d}$, $E_{n}$ is separable; see e.g. [11, page 20]. Let $\{x_{n,i}\}_{i\in\mathbb{N}}$ be an enumeration of a countable dense set in $E_{n}$. Define the ($\mathbb{N}\cup+\infty$)-valued random variable $J_{n}$ to be the first $i$ such that $x_{n,i}\notin Z_{n}({\cal U}^{\prime}_{n,D_{n}})$, or $J_{n}:=+\infty$ if there is no such $i$. Since $E_{n}\setminus Z_{n}({\cal U}^{\prime}_{n,D_{n}})$ is open in $E_{n}$, if $E_{n}\setminus Z_{n}({\cal U}^{\prime}_{n,D_{n}})\neq\varnothing$ then $J_{n}<\infty$. Therefore
$\mathbb{P}[E_{n}\subset Z_{n}({\cal U}_{n,D_{n}})|\{E_{n}\subset Z_{n}({\cal U}^{\prime}_{n,D_{n}})\}^{c}]\leq\mathbb{P}[x_{n,J_{n}}\in Z_{n}({\cal U}^{\prime\prime}_{n,D_{n}})|J_{n}<\infty],$
which tends to zero by (3.11) and the independence of ${\cal U}^{\prime}_{n,D_{n}}$ and ${\cal U}^{\prime\prime}_{n,D_{n}}$, and (3.10) follows. ∎
The following terminology and notation will be used in the proof of Lemma 3.5, and again later on. Given $x\in\mathbb{R}^{d}$, we let $\pi_{1}(x),\ldots,\pi_{d}(x)$ denote the co-ordinates of $x$, and refer to $\pi_{d}(x)$ as the height of $x$. Given ${\bf x}_{1}=(x_{1},s_{1}),\ldots,{\bf x}_{d}=(x_{d},s_{d})\in\mathbb{R}^{d}\times\mathbb{R}_{+}$, if $\cap_{i=1}^{d}\partial B(x_{i},s_{i})$ consists of exactly two points, we refer to these as $p({\bf x}_{1},\ldots,{\bf x}_{d})$ and $q({\bf x}_{1},\ldots,{\bf x}_{d})$ with $p({\bf x}_{1},\ldots,{\bf x}_{d})$ at a smaller height than $q({\bf x}_{1},\ldots,{\bf x}_{d})$ (or if they are at the same height, take $p({\bf x}_{1},\ldots,{\bf x}_{d})<q({\bf x}_{1},\ldots,{\bf x}_{d})$ in the lexicographic ordering). Define the indicator function
$h({\bf x}_{1},\ldots,{\bf x}_{d}):=$ ${\bf 1}\{\pi_{d}(x_{1})\leq\min(\pi_{d}(x_{2}),\ldots,\pi_{d}(x_{d}))\}$
×$\times{\bf 1}\{\#(\cap_{i=1}^{d}\partial B(x_{i},r_{i}))=2\}{\bf 1}\{\pi_{d}(x_{1})<\pi_{d}(q({\bf x}_{1},\ldots,{\bf x}_{d}))\}.$ (3.12)
Given also $n>0$ we also define
$h_{n}({\bf x}_{1},\ldots,{\bf x}_{d}):=h((x_{1},r_{n}s_{1}),\ldots,(x_{d},r_{n}s_{d}));$ (3.13)
$p_{n}({\bf x}_{1},\ldots,{\bf x}_{d}):=p((x_{1},r_{n}s_{1}),\ldots,(x_{d},r_{n}s_{d}));$ (3.14)
$q_{n}({\bf x}_{1},\ldots,{\bf x}_{d}):=q((x_{1},r_{n}s_{1}),\ldots,(x_{d},r_{n}s_{d})).$ (3.15)
Lemma 3.4 (Integrability of $h$).
There exists a constant $c$ depending only on $d$ such that for all $x_{1}\in\mathbb{R}^{d}$ and all $s_{1},\ldots,s_{d}\in\mathbb{R}_{+}$,
[…]
Proof.
Denote the left hand side of (3.16) by $I(x_{1},s_{1},\ldots,s_{d})$. Divide $\mathbb{R}^{d}$ into rectilinear unit hypercubes $Q_{z},z\in\mathbb{Z}^{d}$, with $Q_{z}$ centred at $z$ for each $z$. Then
[…]
$\leq\sum_{z\in\mathbb{Z}^{d}}\int_{\mathbb{R}^{d}}\cdots\int_{\mathbb{R}^{d}}\prod_{i=1}^{d}{\bf 1}\{\partial B(x_{i},s_{i})\cap Q_{z}\neq\varnothing\}dx_{2}\cdots dx_{d}$
[…]
We are now ready to present the main result of this section. In it, $|\Omega|$ denotes the $(d-1)$-dimensional Lebesgue measure of $\Omega$. Recall the definition of $c_{d,1}$ at (2.3).
Lemma 3.5 (Coverage near the boundary of a half-space).
Let $\Omega\subset\mathbb{R}^{d-1}$ and (for each $n>0$) $\Omega_{n}\subset\Omega$ be closed and Riemann measurable. Assume that (2.5) holds for some $\beta\in\mathbb{R}$ or $\beta=+\infty$, and also that $\limsup_{n\to\infty}(nr_{n}^{d}/(\log n))<\infty$. Then
$\lim_{n\to\infty}(\mathbb{P}[\Omega\times\{0\}\subset Z_{n}({\cal U}_{n,\mathbb{H}})])=\exp\left(-c_{d,k}\Big{(}\frac{(\mathbb{E}\,[Y^{d-1}])^{d-1}}{(\mathbb{E}\,[Y^{d}])^{d-2+1/d}}\Big{)}|\Omega|e^{-\beta/2}\right).$ (3.17)
Also, given $\varepsilon>0$ and $a_{n}\in(0,\infty)$ for each $n>0$, such that $a_{n}=O(n^{(1/d)-\varepsilon})$ as $n\to\infty$, we have
$(\displaystyle\lim_{n\to\infty}(\mathbb{P}[\{(\Omega_{n}\times\{0\})\cup($$(\partial\Omega_{n})\times[0,a_{n}r_{n}])\subset Z^{o}_{n}({\cal U}^{\prime}_{n,\mathbb{H}})\}$
$\setminus\{(\Omega_{n}\times[0,a_{n}r_{n}])\subset Z_{n}({\cal U}^{\prime}_{n,\mathbb{H}})\}])=0.$ (3.18)
Remark 3.6.
Usually we shall use Lemma 3.5 in the case where $\beta<\infty$. In this case the extra condition $\limsup_{n\to\infty}(nr_{n}^{d}/(\log n))<\infty$ is automatic. When $\beta=\infty$, in (3.17) we use the convention $e^{-\infty}:=0$.
Proof of Lemma 3.5.
Assume for now that $\beta<\infty$. Considering the slices of balls induced by the points of ${\cal U}_{n,\mathbb{H}}$ that intersect the hyperplane $\mathbb{R}^{d-1}\times\{0\}$, we have a $(d-1)$-dimensional SPBM, where we claim the parameters (in the notation of Lemma 3.2) are
$\delta=r_{n},~{}~{}~{}\lambda=nr_{n}\mathbb{E}\,[Y],~{}~{}~{}~{}\alpha=\omega_{d}\mathbb{E}\,[Y^{d}]/(2\mathbb{E}\,[Y]).$ (3.19)
[…]
$=(\mathbb{E}\,[Y^{d}]/\mathbb{E}\,[Y])\omega_{d}/(2\omega_{d-1}),$ (3.20)
and the asserted value of $\alpha$ follows.
By (3.19), followed by (2.5),
$\alpha\delta^{d-1}\lambda$ $=\Big{(}\frac{\omega_{d}\mathbb{E}\,[Y^{d}]}{2\mathbb{E}\,[Y]}\Big{)}r_{n}^{d-1}nr_{n}\mathbb{E}\,[Y]=(\omega_{d}/2)nr_{n}^{d}\mathbb{E}\,[Y^{d}]$
$=(1-1/d)\log n+(d+k-3+1/d)\log\log n+(\beta/2)+o(1).$
Also by (3.19) and (3.4),
$\log\lambda$ $=(1-1/d)\log n+\Big{(}\frac{1}{d}\Big{)}\log\log n+\log\Big{(}\Big{(}\frac{2-2/d}{\omega_{d}\mathbb{E}\,[Y^{d}]}\Big{)}^{1/d}\mathbb{E}\,[Y]\Big{)}+o(1).$
Also $\log\log\lambda=\log\log n+\log(1-1/d)+o(1).$ Hence
$\alpha\delta^{d-1}\lambda-\log\lambda-(d+k-3)\log\log\lambda=(\beta/2)-\log(c^{\prime\prime}_{d,k,Y})+o(1),$
where we take
$c^{\prime\prime}_{d,k,Y}$ $:=\left(\frac{2-2/d}{\omega_{d}\mathbb{E}\,[Y^{d}]}\right)^{1/d}\mathbb{E}\,[Y](1-1/d)^{d+k-3}=(1-1/d)^{d+k-3+1/d}\left(\frac{2(\mathbb{E}\,[Y])^{d}}{\omega_{d}\mathbb{E}\,[Y^{d}]}\right)^{1/d}.$
By (3.20), and a similar calculation for $\mathbb{E}\,[W^{d-2}]$,
$\frac{(\mathbb{E}\,[W^{d-2}])^{d-1}}{(\mathbb{E}\,[W^{d-1}])^{d-2}}=\left(\frac{\omega_{d-1}\mathbb{E}\,[Y^{d-1}]}{2\omega_{d-2}\mathbb{E}\,[Y]}\right)^{d-1}\left(\frac{\omega_{d}\mathbb{E}\,[Y^{d}]}{2\omega_{d-1}\mathbb{E}\,[Y]}\right)^{2-d}=\frac{\omega_{d-1}^{2d-3}(\mathbb{E}\,[Y^{d-1}])^{d-1}}{2\omega_{d-2}^{d-1}\omega_{d}^{d-2}(\mathbb{E}\,[Y^{d}])^{d-2}\mathbb{E}\,[Y]}.$
Thus by Lemma 3.2 we obtain that
$\lim_{n\to\infty}(\mathbb{P}[\Omega\times\{0\}\subset Z_{n}({\cal U}_{n,\mathbb{H}})])=\exp\Big{(}-\Big{(}\frac{c_{d-1}\omega_{d-1}^{2d-3}(\mathbb{E}\,[Y^{d-1}])^{d-1}}{(k-1)!2\omega_{d-2}^{d-1}\omega_{d}^{d-2}\mathbb{E}\,[Y](\mathbb{E}\,[Y^{d}])^{d-2}}\Big{)}$
×$\times c^{\prime\prime}_{d,k,Y}|\Omega|e^{-\beta/2}\Big{)},$
and hence (3.17).
Having now verified (3.17) in the case where $\beta<\infty$, we can then easily deduce (3.17) in the other case too.
It remains to prove (3.18); we now consider general $\beta\in(0,\infty]$. For the duration of this proof only, let $F_{n}$ be the event displayed in (3.18) but with ${\cal U}^{\prime}_{n,\mathbb{H}}$ changed to ${\cal U}_{n,\mathbb{H}}$ both times, i.e.
$F_{n}:=\{(\Omega_{n}\times\{0\})\cup((\partial\Omega_{n})\times[0,a_{n}r_{n}])\subset Z^{o}_{n}({\cal U}_{n,\mathbb{H}})\}\setminus\{(\Omega_{n}\times[0,a_{n}r_{n}])\subset Z_{n}({\cal U}_{n,\mathbb{H}})\}.$
By Lemma 3.3, to prove (3.18) it suffices to prove that $\mathbb{P}[F_{n}]\to 0$ as $n\to\infty$.
Let $E_{n}$ be the (exceptional) event that there exist $d$ distinct marked points $(x_{1},s_{1}),\ldots,$ $(x_{d},s_{d})$ of ${\cal U}_{n,\mathbb{H}}$ such that $\cap_{i=1}^{d}\partial B(x_{i},r_{n}s_{i})$ has non-empty intersection with the hyperplane $\partial\mathbb{H}$. Then $\mathbb{P}[E_{n}]=0$.
Suppose that $F_{n}\setminus E_{n}$ occurs. Let $w$ be a location of minimal height (i.e., $d$-coordinate) in the closure of $(\Omega_{n}\times[0,a_{n}r_{n}])\setminus Z_{n}({\cal U}_{n,\mathbb{H}})$. Since we assume $(\partial\Omega_{n})\times[0,a_{n}r_{n}]\subset Z^{o}_{n}({\cal U}_{n,\mathbb{H}})$ occurs, $w$ must lie in $\Omega_{n}^{o}\times(0,a_{n}r_{n}]$. Also we claim that $w$ must be a ‘corner’ given by the meeting point of the boundaries of exactly $d$ balls $B(x_{1},r_{n}s_{1}),\ldots,B(x_{d},r_{n}s_{d})$ with $\{(x_{i},s_{i})\}_{i=1}^{d}\subset{\cal U}_{n,\mathbb{H}}$, and with $x_{1}$ the lowest of the $d$ points $x_{1},\ldots,x_{d}$, and with #$\#(\cap_{i=1}^{d}\partial B(x_{i},r_{n}s_{i}))=2$, and $w\notin Z_{n}({\cal U}_{n,\mathbb{H}}\setminus\{(x_{1},s_{1}),\ldots,(x_{d},s_{d})\})$.
Indeed, if $w$ is not at the boundary of any such ball, then for some $\delta>0$ we have $B(w,\delta)\cap Z_{n}({\cal U}_{n,\mathbb{H}})=\varnothing$, and then we could find a location in $(\Omega_{n}\times[0,ar_{n}])\setminus Z_{n}({\cal U}_{n,\mathbb{H}})$ lower than $w$, a contradiction. Next, suppose instead that $w$ lies at the boundary of fewer than $d$ such balls. Then denoting by $L$ the intersection of the supporting hyperplanes at $w$ of each of these balls, we have that $L$ is an affine subspace of $\mathbb{R}^{d}$, of dimension at least 1. Take $\delta>0$ small enough so that $B(w,\delta)$ does not intersect any of the boundaries of balls centred at points of ${\cal U}_{n,\mathbb{H}}$ (viewed here as a marked Poisson point process in $\mathbb{R}^{d}$) and with radius given by their marks times $r_{n}$, other than those which meet at $w$. Taking $w^{\prime}\in L\cap B(w,\delta)\setminus\{w\}$ such that $w^{\prime}$ is at least as low as $w$, we have that $w^{\prime}$ lies in the interior of $Z_{n}({\cal U}_{n,\mathbb{H}})^{c}$. Hence for some $\delta^{\prime}>0$, $B(w^{\prime},\delta^{\prime})\cap Z_{n}({\cal U}_{n,\mathbb{H}})=\varnothing$ and we can find a location in $B(w^{\prime},\delta^{\prime})$ that is lower than $w$, yielding a contradiction for this case too. Finally, with probability 1 there is no set of more than $d$ points of ${\cal U}_{n,\mathbb{H}}$ such that the boundaries of the associated balls have non-empty intersection, so $w$ is not at the boundary of more than $d$ such balls. Thus we have justified the claim.
Moreover $w$ must be the point $q_{n}({\bf x}_{1},\ldots,{\bf x}_{d})$ rather than $p_{n}({\bf x}_{1},\ldots,{\bf x}_{d})$, where for $1\leq i\leq d$ we write ${\bf x}_{i}$ for $(x_{i},r_{i})$, because otherwise by extending the line segment from $q_{n}({\bf x}_{1},\ldots,{\bf x}_{d})$ to $p_{n}({\bf x}_{1},\ldots,{\bf x}_{d})$ slightly beyond $p_{n}({\bf x}_{1},\ldots,{\bf x}_{d})$ we could find a point in $(\Omega_{n}\times[0,ar_{n}])\setminus Z_{n}({\cal U}_{n,\mathbb{H}})$ lower than $w$, contradicting the statement that $w$ is a location of minimal height in the closure of $(\Omega_{n}\times[0,ar_{n}])\setminus Z_{n}({\cal U}_{n,\mathbb{H}})$. Moreover, $w$ must be strictly higher than $x_{1}$, since if $\pi_{d}(w)\leq\min(\pi_{d}(x_{1}),\ldots,\pi_{d}(x_{d}))$, then locations just below $w$ would lie in $(\Omega_{n}\times[0,ar_{n}])\setminus Z_{n}({\cal U}_{n,\mathbb{H}})$, contradicting the statement that $w$ is a point of minimal height in the closure of $(\Omega_{n}\times[0,ar_{n}])\setminus Z_{n}({\cal U}_{n,\mathbb{H}})$. Hence, $h_{n}({\bf x}_{1},\ldots,{\bf x}_{d})=1$, where $h_{n}(\cdot)$ was defined at (3.13).
Thus if $F_{n}\setminus E_{n}$ occurs, then there exists at least one $d$-tuple of marked points ${\bf x}_{1}=(x_{1},s_{1}),\ldots,{\bf x}_{d}=(x_{d},s_{d})\in{\cal U}_{n,\mathbb{H}}$, such that $h_{n}({\bf x}_{1},\ldots,{\bf x}_{d})=1$, and moreover $q_{n}({\bf x}_{1},\ldots,{\bf x}_{d})\in(\Omega_{n}\times(0,a_{n}r_{n}])\setminus Z_{n}({\cal U}_{n,\mathbb{H}}\setminus\{{\bf x}_{1},\ldots,{\bf x}_{d}\})$. Let $M_{n}(a_{n})$ denote the number of such $d$-tuples.
Assume for now that for some constant $a\in(0,\infty)$ we have $a_{n}=a$ for all $n$. We shall prove (3.18) first in this case.
Choose $\delta>0$ such that $\mathbb{P}[Y>\delta]>\delta$. Then choose $c^{\prime}>0$ so that
$\omega_{d-1}\delta^{d}2^{1-d}((\delta/2)\wedge u)\geq c^{\prime}u,~{}~{}~{}\forall u\in[0,a].$
For any $y\in\mathbb{H}$, let $N_{n}(y)$ denote the number of balls making up $Z_{n}({\cal U}_{n,\mathbb{H}})$ which cover $y$, i.e. the number of $(x,s)\in{\cal U}_{n,\mathbb{H}}$ such that $\|x-y\|\leq r_{n}s$. Also, let $\mathbb{H}_{y}:=\{z\in\mathbb{H}:\pi_{d}(z)>\pi_{d}(y)\}$. Then for all $n$ and all $y\in\mathbb{H}$ with $0\leq\pi_{d}(y)\leq ar_{n}$, since for all $s>0$, the half-ball $B(y,r_{n}s)\cap\mathbb{H}_{y}$, and also the cylinder with radius $sr_{n}/2$, upper face centred on $y$ and height $(r_{n}s/2)\wedge\pi_{d}(y)$, are disjoint and contained in $B(y,r_{n}s)\cap\mathbb{H}$,
$\mathbb{E}\,[N_{n}(y)]$ $=n\int_{\mathbb{R}_{+}}|B(y,r_{n}s)\cap\mathbb{H}|\mu_{Y}(ds)$
$\geq n\int_{\mathbb{R}_{+}}(\omega_{d}r_{n}^{d}s^{d}/2)\mu_{Y}(ds)+n\int_{(\delta,\infty)}\omega_{d-1}(r_{n}s/2)^{d-1}((r_{n}s/2)\wedge\pi_{d}(y))\mu_{Y}(ds)$
$\geq(n\omega_{d}/2)r_{n}^{d}\mathbb{E}\,[Y^{d}]+n\delta\omega_{d-1}(\delta r_{n}/2)^{d-1}r_{n}((\delta/2)\wedge(\pi_{d}(y)/r_{n}))$
$\geq(n\omega_{d}/2)r_{n}^{d}\mathbb{E}\,[Y^{d}]+c^{\prime}nr_{n}^{d-1}\pi_{d}(y).$ (3.21)
By the Mecke formula (see e.g. [6]), taking $Y_{1},\ldots,Y_{d}$ to be independent random variables with the distribution of $Y$, there is a constant $c$ such that
$\mathbb{E}\,[M_{n}(a)]\leq$ $cn^{d}\int_{\mathbb{H}}\cdots\int_{\mathbb{H}}\mathbb{E}\,[h_{n}((x_{1},Y_{1}),\ldots,(x_{d},Y_{d}))$
×$\times{\bf 1}\{q_{n}((x_{1},Y_{1}),\ldots,(x_{d},Y_{d}))\in\Omega_{n}\times[0,ar_{n}]\}(k(n\omega_{d}r_{n}^{d})^{k-1})$
×$\times e^{-n(\omega_{d}/2)\mathbb{E}\,[Y^{d}]r_{n}^{d}-c^{\prime}nr_{n}^{d-1}\pi_{d}(q_{n}((x_{1},Y_{1}),\ldots,(x_{d},Y_{d})))}]dx_{d}\cdots dx_{1}.~{}~{}~{}$ (3.22)
Now we change variables to $y_{i}=r_{n}^{-1}(x_{i}-x_{1})$ for $2\leq i\leq d$, noting that
$\pi_{d}(q_{n}((x_{1},Y_{1}),\ldots,(x_{d},Y_{d})))=\pi_{d}(x_{1})+\pi_{d}(q_{n}((o,Y_{1}),(x_{2}-x_{1},Y_{2}),\ldots,(x_{d}-x_{1},Y_{d}))).$
With these changes of variable, we get a factor of $r_{n}^{d(d-1)}$ on changing from $(x_{2},\ldots,x_{d})$ to $(y_{2},\ldots,y_{d})$. Hence by (3.5), there is a constant $c^{\prime\prime}$ such that the expression on the right hand side of (3.22) is at most
[…]
×$\times\int_{\mathbb{H}}\cdots\int_{\mathbb{H}}\mathbb{E}\,[h_{n}((o,Y_{1}),(r_{n}y_{2},Y_{2}),\ldots,(r_{n}y_{d},Y_{d}))e^{-c^{\prime}nr_{n}^{d-1}\pi_{d}(q_{n}((o,Y_{1}),(r_{n}y_{2},Y_{2}),\ldots,(r_{n}y_{d},Y_{d})))}]$
$dy_{d}\ldots dy_{2}.~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}$ (3.23)
In the last expression the first line is bounded by a constant times the expression
$n^{d+k-3+1/d}(r_{n}^{d})^{d+k-2-(d-1)/d}(\log n)^{-(d+k-3+1/d)}=(nr_{n}^{d}/(\log n))^{d+k-3+1/d},$
which is bounded because $nr_{n}^{d}=O(\log n)$ by (3.4). Using Fubini’s theorem and the definitions of $h_{n}(\cdot)$ and $q_{n}(\cdot)$ in (3.13) and (3.15), the expression in the second and third lines of (3.23) can be rewritten as
[…]
$=\int_{\mathbb{R}_{+}^{d}}\int_{\mathbb{H}^{d-1}}h((o,u_{1}),(y_{2},u_{2}),\ldots,(y_{d},u_{d}))e^{-c^{\prime}nr_{n}^{d}\pi_{d}(q((o,u_{1}),(y_{2},u_{2}),\ldots,(y_{d},u_{d})))}$
$d(y_{2},\ldots,y_{d})\mu_{Y}^{d}(d(u_{1},\ldots,u_{d})).$
This tends to zero by dominated convergence because $nr_{n}^{d}\to\infty$ by (2.5) and because the indicator function $({\bf u},{\bf y})\mapsto h((o,u_{1}),(y_{2},u_{2}),\ldots,(y_{d},u_{d}))$ is integrable and is zero when $\pi_{d}(q((o,u_{1}),(y_{2},u_{2}),\ldots,(y_{d},u_{d})))\leq 0$; the integrability follows from Lemma 3.4 and the fact that $\mathbb{E}\,[Y^{d-1}]<\infty$ since we assume $\mathbb{E}\,[Y^{\gamma}]<\infty$ for some $\gamma>d$.
Therefore for any fixed constant $a$ we have
$\lim_{n\to\infty}(\mathbb{E}\,[M_{n}(a)])=0,$ (3.24)
so by Markov’s inequality we obtain that $\mathbb{P}[F_{n}]\leq\mathbb{P}[M_{n}(a)\geq 1]\to 0$, and hence (3.18), in the case where $a_{n}\equiv a$.
Now assuming only that $a_{n}=O(n^{(1/d)-\varepsilon})$ for some $\varepsilon>0$, choose $a>0$ such that
$\int_{(a,\infty)}x^{d}\mu_{Y}(dx)<\varepsilon\mathbb{E}\,[Y^{d}]/4.$
Assume without loss of generality that $a_{n}>a$. For $y\in\mathbb{H}$ with $ar_{n}\leq\pi_{d}(y)\leq a_{n}r_{n}$, instead of (3.21) we use the estimate
$\mathbb{E}\,[N_{n}(y)]$ $=n\int_{\mathbb{R}_{+}}|B(y,r_{n}s)\cap\mathbb{H}|\mu_{Y}(ds)$
$\geq n\int_{(0,a)}(\omega_{d}r_{n}^{d}s^{d})\mu_{Y}(ds)$
$\geq n\omega_{d}r_{n}^{d}\mathbb{E}\,[Y^{d}](1-\varepsilon/4).$
By this, and the Mecke formula, there is a new constant $c$ such that
$\mathbb{E}\,[M_{n}(a_{n})-M_{n}(a)]\leq cn^{d}(nr_{n}^{d})^{k-1}\int_{\mathbb{H}}\cdots\int_{\mathbb{H}}\mathbb{E}\,[h_{n}((x_{1},Y_{1}),\ldots,(x_{d},Y_{d}))$
×$\times{\bf 1}\{q_{n}((x_{1},Y_{1}),\ldots,(x_{d},Y_{d}))\in\Omega_{n}\times[ar_{n},a_{n}r_{n}]\}e^{-n\omega_{d}r_{n}^{d}\mathbb{E}\,[Y^{d}](1-\varepsilon/4)}]dx_{d}\cdots dx_{1}.~{}~{}~{}$ (3.25)
Changing variables as before to $y_{i}=r_{n}^{-1}(x_{i}-x_{1})$ for $2\leq i\leq d$, and using (3.5), we find there is a constant $c^{\prime\prime\prime}$ such that the expression in the right hand side of (3.25) is at most
$c^{\prime\prime\prime}n^{d+k-1}r_{n}^{d(d+k-2)}n^{-(2-2/d)+\varepsilon/2}\mathbb{E}\,\Big{[}\int_{\mathbb{H}}\cdots\int_{\mathbb{H}}h_{n}((o,Y_{1}),(r_{n}y_{2},Y_{2}),\ldots,(r_{n}y_{d},Y_{d}))$
×$\times{\bf 1}_{\Omega_{n}\times[ar_{n},a_{n}r_{n}]}(x_{1}+q_{n}((o,Y_{1}),(r_{n}y_{2},Y_{2}),\ldots,(r_{n}y_{d},Y_{d})))dx_{1}dy_{d}\ldots dy_{2}\Big{]}$
$\leq c^{\prime\prime\prime}(nr_{n}^{d})^{d+k-2}n^{-1+(2/d)+\varepsilon/2}|\Omega_{n}|(a_{n}-a)r_{n}$
×$\times\mathbb{E}\,\Big{[}\int_{\mathbb{H}^{d-1}}h((o,Y_{1}),(y_{2},Y_{2}),\ldots,(y_{d},Y_{d}))d((y_{2},\ldots,y_{d}))\Big{]}.$
Since $\mathbb{E}\,[Y^{d-1}]<\infty$ the expectation in the last line is finite by Lemma 3.4, and since also $a_{n}=O(n^{(1/d)-\varepsilon})$ we obtain that
$\mathbb{E}\,[M_{n}(a_{n})-M_{n}(a)]=O((nr_{n}^{d})^{d+k-2+1/d}n^{-1+(2/d)+(\varepsilon/2)-\varepsilon}),$
which tends to zero. Combined with (3.24) this shows that $\mathbb{E}\,[M_{n}(a_{n})]\to 0$, and therefore by Markov’s inequality we obtain that $\mathbb{P}[F_{n}]\leq\mathbb{P}[M_{n}(a_{n})\geq 1]\to 0$, and hence (3.18). ∎
3.3 Proof of Theorem 2.1
In this subsection, we set $d=2$ and take $A$ to be polygonal. Denote the vertices of $A$ by $q_{1},\ldots,q_{\kappa}$, and the angles subtended at these vertices by $\alpha_{1},\ldots,\alpha_{\kappa}$ respectively.
For the duration of this subsection (and the next) we fix $\beta\in\mathbb{R}$. Assume we are given real numbers $(r_{n})_{n>0}$ satisfying the case $d=2$ of (2.5), i.e.
$\lim_{n\to\infty}\left(\pi nr_{n}^{2}\mathbb{E}\,[Y^{2}]-\log n-(2k-1)\log\log n\right)=\beta.$ (3.26)
Fix ${\zeta}>0$. For $n>0$, in this subsection we define the ‘corner regions’ $Q_{n}$ and $Q_{n}^{-}$ by
$Q_{n}:=\cup_{j=1}^{\kappa}B(q_{j},n^{3{\zeta}}r_{n})\cap A;~{}~{}~{}~{}~{}Q^{-}_{n}:=\cup_{j=1}^{\kappa}B(q_{j},n^{2{\zeta}}r_{n})\cap A.$
Also we define ${\cal U}_{n,A}$, $Z_{n}({\cal U}_{n,A})$ and $Z^{o}_{n}({\cal U}_{n,A})$ by (3.6), (3.7) and (3.8).
Lemma 3.7 (Coverage of regions near the corners of $A$).
Provided ${\zeta}$ is taken to be small enough, $\mathbb{P}[Q_{n}\subset Z_{n}^{o}({\cal U}_{n,A})]\to 1$ as $n\to\infty$.
Proof.
Let $\delta>0$ with $\mathbb{P}[Y>3\delta]>\delta$. Then the restricted SPBM $Z^{o}_{n}({\cal U}_{n,A})$ stochastically dominates a Poisson Boolean model $Z^{\prime}_{n}$ with closed balls of deterministic radius $2\delta r_{n}$, centred on the points of a homogeneous Poisson point process in $A$ of intensity $\delta n$. Let $Z^{\prime\prime}_{n}$ be a Poisson Boolean model with closed balls of deterministic radius $\delta r_{n}$, centred on the points of a homogeneous Poisson point process in $A$ of intensity $\delta n$.
There is a constant $K$, independent of $n$, such that for all $n\geq 1$ there exist points $y_{n,1},\ldots,y_{n,\lfloor Kn^{6{\zeta}}\rfloor}\in A$ such that $Q_{n}\subset\cup_{j=1}^{\lfloor Kn^{6{\zeta}}\rfloor}B(y_{n,j},\delta r_{n})$. Also there is a constant $a>0$ such that for all $n\geq 1$ and all $y\in A$ we have $|B(y,\delta r_{n})\cap A|\geq ar_{n}^{2}$. Then
$\mathbb{P}[Q_{n}\setminus Z_{n}^{o}({\cal U}_{n,A})\neq\varnothing]$ $\leq\mathbb{P}[Q_{n}\setminus Z^{\prime}_{n}\neq\varnothing]$
$\leq\mathbb{P}[\cup_{j=1}^{\lfloor Kn^{6{\zeta}}\rfloor}\{y_{n,j}\notin Z^{\prime\prime}_{n}\}]$
$\leq Kn^{6{\zeta}}k(n\pi r_{n}^{2})^{k-1}\exp(-a\delta nr_{n}^{2}),$ (3.27)
and since $(nr_{n}^{2})/\log n\to 1/(\pi\mathbb{E}\,[Y^{2}])$ by (3.26), provided we take ${\zeta}<a\delta/(6\pi\mathbb{E}\,[Y^{2}])$ the expression in (3.27) tends to zero, and the result follows. ∎
Recall from (3.6) the definition of the point processes ${\cal U}^{\prime}_{n,D}$ for $n>0,D\subset\mathbb{R}^{d}$.
Lemma 3.8 (Coverage of regions near the edges of $A$).
Assume ${\zeta}<1/d$. Then
$\lim_{n\to\infty}(\mathbb{P}[(\partial A\setminus Q^{-}_{n})\subset Z_{n}({\cal U}^{\prime}_{n,A})])=\exp\Big{(}-\Big{(}\frac{c_{2,k}\mathbb{E}\,[Y]}{(\mathbb{E}\,[Y^{2}])^{1/2}}\Big{)}|\partial A|e^{-\beta/2}\Big{)}.$ (3.28)
Also,
$\lim_{n\to\infty}(\mathbb{P}[\{A^{(3n^{\zeta}r_{n})}\subset Z_{n}({\cal U}^{\prime}_{n,A})\}\cap\{(\partial A\cup Q_{n})\subset Z^{o}_{n}({\cal U}^{\prime}_{n,A})\}\setminus\{A\subset Z_{n}({\cal U}^{\prime}_{n,A})\}])=0.$ (3.29)
Proof.
Denote the line segments making up $\partial A$ by $I_{1},\ldots,I_{\kappa}$, and for $n>0$ and $1\leq i\leq\kappa$ set $I_{n,i}:=I_{i}\setminus Q^{-}_{n}$.
Let $i,j,k\in\{1,\ldots,\kappa\}$ be such that $i\neq j$ and the edges $I_{i}$ and $I_{j}$ are both incident to $q_{k}$. Provided $n$ is large enough, if $x\in I_{n,i}$ and $y\in I_{n,j}$, then $\|x-y\|\geq(n^{2\zeta}r_{n})\sin\alpha_{k}\geq 3n^{\zeta}r_{n}$. Hence for all large enough $n$ the events $\{I_{n,1}\subset Z_{n}({\cal U}^{\prime}_{n,A})\},\ldots,\{I_{n,\kappa}\subset Z_{n}({\cal U}^{\prime}_{n,A})\}$ are mutually independent. Therefore
$\mathbb{P}[(\partial A\setminus Q^{-}_{n})\subset Z_{n}({\cal U}^{\prime}_{n,A})]=\prod_{i=1}^{\kappa}\mathbb{P}[I_{n,i}\subset Z_{n}({\cal U}^{\prime}_{n,A})],$
and by Lemmas 3.5 and 3.3, this converges to the right hand side of (3.28).
Now we prove (3.29). For $n>0$, and $i\in\{1,2,\ldots,\kappa\}$, let $S_{n,i}$ denote the rectangular block of dimensions $|I_{n,i}|\times 3n^{\zeta}r_{n}$, consisting of all points in $A$ at perpendicular distance at most $3n^{\zeta}r_{n}$ from $I_{n,i}$. Let $\partial_{{\rm side}}S_{n,i}$ denote the union of the two ‘short’ edges of $S_{n,i}$, i.e. the two edges bounding $S_{n,i}$ which are perpendicular to $I_{n,i}$.
Then for $n$ large, $A\setminus(A^{(3n^{\zeta}r_{n})}\cup Q_{n})\subset\cup_{i=1}^{\kappa}S_{n,i}$, and also $\partial_{\rm side}S_{n,i}\subset Q_{n}$ for $1\leq i\leq\kappa$, so that
$\{(A^{(3n^{\zeta}r_{n})}\cup\partial A)\subset Z_{n}({\cal U}^{\prime}_{n,A})\}\cap\{Q_{n}\subset Z^{o}_{n}({\cal U}^{\prime}_{n,A})\}\setminus\{A\subset Z_{n}({\cal U}^{\prime}_{n,A})\}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}$
$\subset\cup_{i=1}^{\kappa}[\{I_{n,i}\subset Z_{n}({\cal U}^{\prime}_{n,A})\}\cup\{\partial_{{\rm side}}S_{n,i}\subset Z_{n}^{o}({\cal U}^{\prime}_{n,A})\}\setminus\{S_{n,i}\subset Z_{n}({\cal U}^{\prime}_{n,A})\}].$
For $i\in\{1,\ldots,\kappa\}$, let $I^{\prime}_{n,i}$ denote an interval of length $|I_{n,i}|$ in $\mathbb{R}$. Using the Mapping theorem [6] and the translation and rotation invariance of Lebesgue measure one sees that
[$\mathbb{P}[\{I_{n,i}\subset Z_{n}({\cal U}^{\prime}_{n,A})\}\cap\{\partial_{{\rm side}}S_{n,i}\subset Z^{o}_{n}({\cal U}^{\prime}_{n,A})\}\setminus\{S_{n,i}\subset Z_{n}({\cal U}^{\prime}_{n,A})\}]$
=[$=\mathbb{P}[\{(I^{\prime}_{n,i}\times\{0\})\subset Z_{n}({\cal U}^{\prime}_{n,\mathbb{H}})\}\cap\{(\partial I^{\prime}_{n,i})\times[0,3n^{\zeta}r_{n}]\subset Z_{n}^{o}({\cal U}^{\prime}_{n,\mathbb{H}})\}$
$\setminus\{(I^{\prime}_{n,i}\times[0,3n^{\zeta}r_{n}])\subset Z_{n}({\cal U}^{\prime}_{n,\mathbb{H}})\}].$
By (3.18) from Lemma 3.5, this probability tends to zero, and hence (3.29). ∎
Proof of Theorem 2.1.
Suppose $(r_{n})_{n>0}$ satisfies (3.26). We assert that provided ${\zeta}<1/2$,
$\lim_{n\to\infty}\mathbb{P}[A^{(3n^{\zeta}r_{n})}\subset Z_{n}({\cal U}^{\prime}_{n,A})]=\exp\Big{(}-{\bf 1}_{\{k=1\}}\Big{(}\frac{(\mathbb{E}\,[Y])^{2}}{\mathbb{E}\,[Y^{2}]}\Big{)}|A|e^{-\beta}\Big{)}.$ (3.30)
Indeed, setting $\lambda=n$ and $\delta(\lambda)=r_{n}$, we have that if $k=1$ then (3.2) holds, while if $k\geq 2$ then the left hand side of (3.2) tends to +$\infty$+ ∞. Also $\mathbb{P}[A^{(3n^{\zeta}r_{n})}\subset Z_{n}({\cal U}^{\prime}_{n,A})]$ is bounded from below by $\mathbb{P}[A\subset Z_{n}({\cal U}^{\prime}_{n,\mathbb{R}^{2}})]$, which converges to the right hand side of (3.30) by Lemmas 3.2 and 3.3. Also, if $k=1$, then given $\varepsilon>0$, for $n$ large the event $\{A^{(3n^{\zeta}r_{n})}\subset Z_{n}({\cal U}^{\prime}_{n,A})\}$ is contained in $\{A^{[\varepsilon]}\subset Z_{n}({\cal U}^{\prime}_{n,A})\}$, so by Lemmas 3.2 and 3.3 again
$\limsup_{n\to\infty}\mathbb{P}[A^{(n^{\zeta}r_{n})}\subset Z_{n}({\cal U}^{\prime}_{n,A})]$ $\leq\limsup_{n\to\infty}\mathbb{P}[A^{[\varepsilon]}\subset Z_{n}({\cal U}^{\prime}_{n,A})]$
$=\exp\Big{(}-\Big{(}\frac{(\mathbb{E}\,[Y])^{2}}{\mathbb{E}\,[Y^{2}]}\Big{)}|A^{[\varepsilon]}|e^{-\beta}\Big{)},$
and since $|A^{[\varepsilon]}|\to|A|$ as $\varepsilon\downarrow 0$, this gives us the assertion (3.30).
Also, by (3.28) and Lemma 3.7,
$\lim_{n\to\infty}(\mathbb{P}[\partial A\subset Z_{n}({\cal U}^{\prime}_{n,A})])=\exp\Big{(}-\Big{(}\frac{c_{2,k}\mathbb{E}\,[Y]}{\sqrt{\mathbb{E}\,[Y^{2}]}}\Big{)}|\partial A|e^{-\beta/2}\Big{)}.$ (3.31)
Note [$\mathbb{P}[\{(Q_{n}\cup\partial A)\subset Z_{n}({\cal U}^{\prime}_{n,A})\}\setminus\{(Q_{n}\cup\partial A)\subset Z^{o}_{n}({\cal U}^{\prime}_{n,A})\}]=0$. Therefore using (3.29) followed by Lemma 3.7, and then Lemma 3.3, we obtain that
$\lim_{n\to\infty}\mathbb{P}[A\subset Z_{n}({\cal U}^{\prime}_{n,A})]$ $=\lim_{n\to\infty}\mathbb{P}[(A^{(3n^{\zeta}r_{n})}\cup\partial A\cup Q_{n})\subset Z_{n}({\cal U}^{\prime}_{n,A})]$
$=\lim_{n\to\infty}\mathbb{P}[(A^{(3n^{\zeta}r_{n})}\cup\partial A)\subset Z_{n}({\cal U}^{\prime}_{n,A})],$ (3.32)
provided these limits exist.
The events $\{\partial A\subset Z_{n}({\cal U}^{\prime}_{n,A})\}$ and $\{A^{(3n^{\zeta}r_{n})}\subset Z_{n}({\cal U}^{\prime}_{n,A})\}$ are independent since the first of these events is determined by the configuration of Poisson points with projection onto $\mathbb{R}^{2}$ distant at most $n^{\zeta}r_{n}$ from $\partial A$, while the second event is determined by the Poisson points with projection distant at least $2n^{\zeta}r_{n}$ from $\partial A$. Therefore the limit in (3.32) does indeed exist, and is the product of the limits arising in (3.30) and (3.31). By Lemma 3.3 we get the same limit for $\mathbb{P}[A\subset Z_{n}({\cal U}_{n,A})]$ as for $\mathbb{P}[A\subset Z_{n}({\cal U}^{\prime}_{n,A})]$. This gives us (2.1). ∎
3.4 First steps toward proving Theorem 2.2
In this subsection we assume that $d\geq 2$ and $\partial A\in C^{2}$. Let $k\in\mathbb{N},\beta\in\mathbb{R}$, and assume that $(r_{n})_{n>0}$ satisfies (2.5) and hence also (3.4) and (3.5). Given any $n>0$, Borel $D\subset\mathbb{R}^{d}$, and any point process $\mathcal{X}$ in $\mathbb{R}^{d}\times\mathbb{R}_{+}$, define $F_{n}(D,\mathcal{X})$ to be the event that $D$ is ‘fully $k$-covered’ by the SPBM based on $\mathcal{X}$ with radii scaled by $r_{n}$, that is,
$F_{n}(D,\mathcal{X}):=\{D\subset Z_{n}(\mathcal{X})\},$ (3.33)
where $Z_{n}(\cdot)$ was defined at (3.7). Also, given $r>0$, define the ‘covering number’
$\kappa(D,r):=\min\{m\in\mathbb{N}:\exists x_{1},\ldots,x_{m}\in D~{}{\rm with}~{}D\subset\cup_{i=1}^{m}B(x_{i},r)\}.$ (3.34)
Given $x\in\partial A$ we can express $\partial A$ locally in a neighbourhood of $x$, after a rotation, as the graph of a $C^{2}$ function from $\mathbb{R}^{d-1}$ to $\mathbb{R}$ with all of its partial derivatives taking the value zero at $x$. We shall approximate to that function by the graph of a piecewise affine function (in $d=2$, a piecewise linear function).
[…]
By a compactness argument, we can and do take a finite collection of points $x_{1},\ldots,x_{J}\in\partial A$ such that
$\partial A\subset\cup_{j=1}^{J}B(x_{j},r(x_{j})).$ (3.35)
Then there are constants $\varepsilon_{j}>0$, and rigid motions $\theta_{j},1\leq j\leq J$, such that for each $j$ the set $\theta_{j}(\partial A\cap{\cal N}_{x_{j}})$ is the graph of a $C^{2}$ function $f_{j}$ defined on a ball $I_{j}$ in $\mathbb{R}^{d-1}$, with $\langle f^{\prime}_{j}(u),e\rangle\leq 1/9$ for all $u\in I_{j}$ and all unit vectors $e\in\mathbb{R}^{d-1}$, and also with $\varepsilon_{j}\leq f_{j}(u)\leq 2\varepsilon_{j}$ for all $u\in I_{j}$ and $\theta_{j}(A)\cap(I_{j}\times[0,3\varepsilon_{j}])=\{(u,z):u\in I_{j},0\leq z\leq f(u)\}$.
Later we shall refer to each tuple $(x_{j},r(x_{j}),\theta_{j},f_{j})$, $1\leq j\leq J$, as a chart. More loosely we shall also refer to each set $B(x_{j},r(x_{j}))$ as a chart.
Let $\Gamma\subset\partial A$ be a closed set such that $\Gamma\subset B(x_{j},r(x_{j}))$ for some $j\in\{1,\ldots,J\}$, and such that $\kappa(\partial\Gamma,r)=O(r^{2-d})$ as $r\downarrow 0$, where we set $\partial\Gamma:=\Gamma\cap\overline{\partial A\setminus\Gamma}$, the boundary of $\Gamma$ relative to $\partial A$. To simplify notation we shall assume that $\Gamma\subset B(x_{1},r(x_{1}))$, and moreover that $\theta_{1}$ is the identity map. Then $\Gamma=\{(u,f_{1}(u)):u\in U\}$ for some bounded set $U\subset\mathbb{R}^{d-1}$. Also, writing $\phi(\cdot)$ for $f_{1}(\cdot)$ from now on, we assume
$\phi(U)\subset[\varepsilon_{1},2\varepsilon_{1}]$ (3.36)
and
$A\cap(U\times[0,3\varepsilon_{1}])=\{(u,z):u\in U,0\leq z\leq\phi(u)\}.$ (3.37)
Note that for any $u,v\in U$, by the Mean Value theorem we have for some $w\in[u,v]$ that
$|\phi(v)-\phi(u)|=|\langle v-u,\phi^{\prime}(w)\rangle|\leq(1/9)\|v-u\|.$ (3.38)
Choose (and keep fixed for the rest of this paper) a constant ${\zeta}$ with
$0<{\zeta}<1/(198d(18+d)).$ (3.39)
Henceforth we shall use this choice of ${\zeta}$ in the definition of ${\cal U}^{\prime}_{n,D}$ at (3.9). We shall use $n^{c{\zeta}}r_{n}$, for various choices of constant $c$, to provide various different length scales.
We shall approximate to $\Gamma$ by a polytopal surface $\Gamma_{n}$ with face diameters that are $\Theta(n^{9{\zeta}}r_{n})$, taking all the faces of $\Gamma_{n}$ to be $(d-1)$-dimensional simplices. Later we shall fit together a finite number of surfaces like $\Gamma$ to make up the whole of $\partial A$.
For the polytopal approximation, divide $\mathbb{R}^{d-1}$ into hypercubes of dimension $d-1$ and side $n^{9{\zeta}}r_{n}$, and divide each of these hypercubes into $(d-1)!$ simplices (we take these simplices to be closed). Let $U_{n}$ be the union of all those simplices in the resulting tessellation of $\mathbb{R}^{d-1}$ into simplices, that are contained within $U$, and let $U_{n}^{-}$ be the union of those simplices in the tessellation which are contained within $U^{(n^{10{\zeta}}r_{n})}$, where for $r>0$ we set $U^{(r)}$ to be the set of $x\in U$ at a Euclidean distance more than $r$ from $\mathbb{R}^{d-1}\setminus U$. If $d=2$, the simplices are just intervals. See Figure 2.
Figure 2: Example in $d=3$. The outer crescent-shaped region is $U$, while the inner crescent is $U^{(10{\zeta}r_{n})}$ The outer polygon is $U_{n}$, while the inner polygon is $U_{n}^{-}$
Let $\sigma^{-}(n)$ denote the number of simplices making up $U_{n}^{-}$. Choose $n_{0}>0$ such that $\sigma^{-}(n)>0$ for all $n\geq n_{0}$.
Let $\psi_{n}:U_{n}\to\mathbb{R}$ be the function that is affine on each of the simplices making up $U_{n}$, and agrees with the function $\phi$ on each of the vertices of these simplices. Our approximating surface (or polygonal line if $d=2$) will be defined by $\Gamma_{n}:=\{(u,\psi_{n}(u)-Kn^{-9{\zeta}}r_{n}):u\in U^{-}_{n}\}$, with the constant $K$ given by the following lemma. This lemma uses Taylor expansion to show that $\psi_{n}$ a good approximation to $\phi$.
Lemma 3.9 (Polytopal approximation).
Set $K:=\sup_{n\geq n_{0},u\in U_{n}}((n^{9{\zeta}}r_{n})^{-2}|\phi(u)-\psi_{n}(u)|)$. Then $K<\infty$.
Proof.
See [10, Lemma 7.18]. The notation $t$ there is equivalent to $n$ here. the length scale of $n^{9\xi}r_{n}$ here plays the role of $t^{-\gamma}$ there. ∎
We now subtract a constant from $\psi_{n}$ to obtain a piecewise affine function $\phi_{n}$ that approximates $\phi$ from below. For $n\geq n_{0}$ and $u\in U_{n}$, define $\phi_{n}(u):=\psi_{n}(u)-Kn^{18{\zeta}}r_{n}^{2}$, with $K$ given by Lemma 3.9. Then for all $n\geq n_{0},u\in U_{n}$ we have $|\psi_{n}(u)-\phi(u)|\leq Kn^{18{\zeta}}r_{n}^{2}$ so that
$\phi_{n}(u)\leq\phi(u)\leq\phi_{n}(u)+2Kn^{18{\zeta}}r_{n}^{2}.$ (3.40)
Define the set $\Gamma_{n}:=\{(u,\phi_{n}(u)):u\in U_{n}^{-}\}$. We refer to each $(d-1)$-dimensional face of $\Gamma_{n}$ (given by the graph of $\phi_{n}$ restricted to one of the simplices in our triangulation of $\mathbb{R}^{d-1}$) as simply a face of $\Gamma_{n}$. Denote these faces of $\Gamma_{n}$ by $H_{n,1},\ldots,H_{n,\sigma^{-}(n)}$. The number of faces, $\sigma^{-}(n)$, is $\Theta((n^{9{\zeta}}r_{n})^{1-d})$ as $n\to\infty$. The perimeter (i.e., the $(d-2)$-dimensional Hausdorff measure of the boundary) of each individual face is $\Theta((n^{9{\zeta}}r_{n})^{d-2})$. For $1\leq i\leq\sigma^{-}(n)$, let $\partial_{d-2}H_{n,i}$ denote the boundary of $H_{n,i}$, which is the image under the mapping $u\mapsto(u,\phi_{n}(u))$, of the boundary of the simplex in $\mathbb{R}^{d-1}$ that is the pre-image of $H_{n,i}$ under that mapping.
For $n\geq n_{0}$, define subsets $A_{n},A_{n}^{-},W_{n},A_{n}^{*},W_{n}^{*}$ of $\mathbb{R}^{d}$ (illustrated in Figure 3) by
$A_{n}:=\{(u,z):u\in U_{n},0\leq z\leq\phi(u)\},~{}~{}~{}W_{n}:=\{(u,z):u\in U_{n},0\leq z\leq\phi_{n}(u)\},$
$A_{n}^{-}:=\{(u,z):u\in U_{n}^{-},0\leq z\leq\phi(u)\},~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}$
$A_{n}^{*}:=\{(u,z):u\in U_{n}^{-},\phi_{n}(u)-(3/2)n^{\zeta}r_{n}\leq z\leq\phi(u)\},~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}$
$W_{n}^{*}:=\{(u,z):u\in U_{n}^{-},\phi_{n}(u)-(3/2)n^{\zeta}r_{n}\leq z\leq\phi_{n}(u)\}.~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}$ (3.41)
Thus $A_{n}$ is a ‘thick slice’ of $A$ near the boundary region $\Gamma$, $W_{n}$ is an approximating region having $\Gamma_{n}$ as its upper boundary, and $A_{n}^{*}$, $W_{n}^{*}$ are ‘thin slices’ of $A$ also having $\Gamma$, respectively $\Gamma_{n}$, as upper boundary. By (3.40), (3.36) and (3.37), $W_{n}^{*}\subset A_{n}^{*}\subset A_{n}^{-}\subset A_{n}\subset A$, and $W_{n}^{*}\subset W_{n}\subset A_{n}$.
[…]
Figure 3: When $d=2$ the sets $A_{n},W_{n},A_{n}^{-},A_{n}^{*},W_{n}^{*}$ are approximately rectangular but with a curved upper boundary for $A_{n},A_{n}^{-}$ and $A_{n}^{*}$, a polygonal upper boundary for $W_{n}^{*}$ and $W_{n}$, and a polygonal lower boundary for $A_{n}^{*}$ and $W_{n}^{*}$.
The rest of this subsection, and the next subsection, are devoted to proving the following intermediate step towards a proof of Theorem 2.10. Recall the definition of $c_{d,k}$ at (2.3), and define
$c_{d,k,Y}:=\frac{c_{d,k}(\mathbb{E}\,[Y^{d-1}])^{d-1}}{(\mathbb{E}\,[Y^{d}])^{d-2+1/d}}.$ (3.42)
Proposition 3.10 (Limiting coverage probability for approximating polytopal shell).
It is the case that $\lim_{n\to\infty}\mathbb{P}[F_{n}(W_{n}^{*},{\cal U}^{\prime}_{n,W_{n}})]=\exp(-c_{d,k,Y}|\Gamma|e^{-\beta/2}),$ where $|\Gamma|$ denotes the $(d-1)$-dimensional Hausdorff measure of $\Gamma$.
The following corollary of Lemma 3.9 is a first step towards proving this.
Lemma 3.11.
(a) It is the case that $|A_{n}\setminus W_{n}|=O(n^{18{\zeta}}r_{n}^{2})$ as $n\to\infty$.
(b) Let $K$ be as given in Lemma 3.9, and fix $L>0$. Then for all $n\geq n_{0}$ and $x\in U_{n}^{(Lr_{n})}\times\mathbb{R}$, $|B(x,Lr_{n})\cap A_{n}\setminus W_{n}|\leq 2K\omega_{d-1}L^{d-1}n^{18{\zeta}}r_{n}^{d+1}$.
Proof.
[…]
For (b), let $x\in U_{n}^{(Lr_{n})}\times\mathbb{R}$, and let $u\in U_{n}^{(Lr_{n})}$ be the projection of $x$ onto the first $d-1$ coordinates. Then if $y\in B(x,Lr_{n})\cap A_{n}\setminus W_{n}$, we have $y=(v,s)$ with $\|v-u\|\leq Lr_{n}$ and $\phi_{n}(v)<s\leq\phi(v)$. Therefore using (3.40) yields
[…]
where the integral is a $(d-1)$-dimensional Lebesgue integral. This gives part (b). ∎
The next lemma says that small balls centred in $A_{n}^{*}$ have almost half of their volume in $W_{n}$.
Lemma 3.12.
Let $\varepsilon\in(0,1)$. Then for all large enough $n$, all $x\in A_{n}^{*}$, and all $s\in[\varepsilon r_{n},r_{n}/\varepsilon]$, we have $|B(x,s)\cap W_{n}|>(1-\varepsilon)(\omega_{d}/2)s^{d}$.
Proof.
For all large enough $n$, all $x\in A_{n}^{*}$ and $s\in[\varepsilon r_{n},r_{n}/\varepsilon]$, we have $B(x,s)\cap A\subset A_{n}$, so $B(x,s)\cap A_{n}=B(x,s)\cap A$, and hence by Lemma 3.1 and Lemma 3.11(b),
$|B(x,s)\cap W_{n}|$ $=$ $|B(x,s)\cap A_{n}|-|B(x,s)\cap A_{n}\setminus W_{n}|$
$\geq$ $(1-\varepsilon/2)(\omega_{d}/2)s^{d}-O(n^{18{\zeta}}r_{n}^{d+1}).$
Since $n^{18{\zeta}}r_{n}\to 0$ by (2.5) and (3.39), this gives us the result. ∎
Recall that $\mathbb{H}$ and ${\cal U}^{\prime}_{n,D}$ were defined in Section 3.1 and at (3.9), respectively. The next lemma provides a bound on the probability that a region of diameter $O(r_{n})$ within $A$ or $A_{n}^{*}$ is not fully covered. This will be used for dealing with ‘exceptional’ regions such as those near the boundaries of faces in the polytopal approximation.
Lemma 3.13.
Let $\varepsilon\in(0,1)$, $K_{1}>0$. Then as $n\to\infty$,
$\sup_{z\in\mathbb{R}^{d}}\mathbb{P}[F_{n}(B(z,K_{1}r_{n})\cap A_{n}^{*},{\cal U}^{\prime}_{n,W_{n}})^{c}]=O(n^{\varepsilon-(d-1)/d}),$ (3.43)
$\sup_{z\in\mathbb{R}^{d}}\mathbb{P}[F_{n}(B(z,K_{1}r_{n})\cap\mathbb{H},{\cal U}^{\prime}_{n,\mathbb{H}})^{c}]=O(n^{\varepsilon-(d-1)/d}),$ (3.44)
and
$\sup_{z\in\mathbb{R}^{d}}\mathbb{P}[F_{n}(B(z,K_{1}r_{n})\cap A,{\cal U}^{\prime}_{n,A})^{c}]=O(n^{\varepsilon-(d-1)/d}).$ (3.45)
Proof.
Let $\delta\in(0,1/2)$ with $(1-\delta)^{d+1}\mathbb{E}\,[Y^{d}{\bf 1}_{\{\delta\leq Y\leq 1/\delta\}}]>(1-\varepsilon/2)\mathbb{E}\,[Y^{d}]$. For any point set $\mathcal{X}\subset\mathbb{R}^{d}\times[\delta,\infty)$, define $\tilde{Z}_{n}(\mathcal{X})$ similarly to $Z_{n}(\mathcal{X})$ but with slightly smaller balls, namely
$\tilde{Z}_{n}(\mathcal{X}):=\{y\in\mathbb{R}^{d}:\#(\{(x,s)\in\mathcal{X}:y\in B(x,(1-\delta)r_{n}s)\})\geq k\},$
and note that for any $y\in\tilde{Z}_{n}(\mathcal{X})$ we have $B(y,\delta^{2}r_{n})\subset Z_{n}(\mathcal{X})$.
There is a constant $\ell$ independent of $n$ such that given $z\in\mathbb{R}^{d}$ and given $n$, we can (and do) choose $x_{n,1},\ldots,x_{n,\ell}\in A_{n}^{*}$ with $B(z,K_{1}r_{n})\cap A^{*}_{n}\subset\cup_{i=1}^{\ell}B(x_{n,i},\delta^{2}r_{n})$. Then for all $n$ large enough, and for $1\leq i\leq\ell$, using Lemma 3.12 in the third line below we have
$1-\mathbb{P}[F_{n}(B(x_{n,i},\delta^{2}r_{n})\cap A^{*}_{n},{\cal U}^{\prime}_{n,W_{n}})]\leq\mathbb{P}[x_{n,i}\notin\tilde{Z}_{n}({\cal U}^{\prime}_{n,W_{n}}\cap(W_{n}\times[\delta,1/\delta]))]$
$\leq k(n\omega_{d}(r_{n}/\delta)^{d})^{k-1}\exp\Big{(}-n\int_{[\delta,1/\delta]}|W_{n}\cap B(x_{i},(1-\delta)r_{n}t)|\mu_{Y}(dt)\Big{)}$
$\leq k(n\omega_{d}(r_{n}/\delta)^{d})^{k-1}\exp\Big{(}-n\int_{[\delta,1/\delta]}(1-\delta)^{d+1}(\omega_{d}/2)r_{n}^{d}t^{d}\mu_{Y}(dt)\Big{)}$
$\leq k(n\omega_{d}(r_{n}/\delta)^{d})^{k-1}\exp\Big{(}-(\omega_{d}/2)nr_{n}^{d}(1-\varepsilon/2)\mathbb{E}\,[Y^{d}]\Big{)}.$
Therefore by(3.5), for $n$ large
[$\mathbb{P}[\{F_{n}(B(x_{n,i},\delta^{2}r_{n})\cap A_{n}^{*},{\cal U}^{\prime}_{n,W_{n}})\}^{c}]\leq n^{\varepsilon-(d-1)/d}.$
Taking the union of the above events for $1\leq i\leq\ell$, and applying the union bound, gives us (3.43). The proofs of (3.44) and (3.45) are similar. ∎
Let $\partial_{d-2}\Gamma_{n}:=\cup_{i=1}^{\sigma^{-}(t)}\partial_{d-2}H_{n,i}$, the union of all $(d-2)$-dimensional faces in the boundaries of the faces making up $\Gamma_{n}$ (the $H_{n,i}$ were defined just after (3.40)). Recall from (3.39) that ${\zeta}\in(0,1/(198d(18+d)))$. Given $n>0$, define the set $Q^{+}_{n}\subset\mathbb{R}^{d}$ by
$Q^{+}_{n}:=(\partial_{d-2}\Gamma_{n}\oplus B(o,8dn^{4{\zeta}}r_{n}))\cap A_{n}^{*}.$ (3.46)
Thus $Q_{n}^{+}$ is a region near the corners of our polygon approximating $\partial A$ (if $d=2$) or near the boundaries of the faces of our polytopal surface approximating $\partial A$ (if $d\geq 3$). In the next lemma we show that $Q_{n}^{+}$ is fully $k$-covered with high probability.
Lemma 3.14.
It is the case that $\mathbb{P}[F_{n}(Q^{+}_{n},{\cal U}^{\prime}_{n,W_{n}})]\to 1$ as $n\to\infty$.
Proof.
Let $\varepsilon:={\zeta}/2$. For each face $H_{n,i}$ of $\Gamma_{n}$, $1\leq i\leq\sigma^{-}(n)$, we claim that we can take $x_{i,1},\ldots,x_{i,k_{n,i}}\in\mathbb{R}^{d}$ with $\max_{1\leq i\leq\sigma^{-}(n)}k_{n,i}=O(n^{9d{\zeta}-10{\zeta}})$, such that
$(\partial_{d-2}H_{n,i})\oplus B(o,8dn^{4{\zeta}}r_{n})\subset\cup_{j=1}^{k_{n,i}}B(x_{i,j},r_{n}).$ (3.47)
Indeed, we can cover $\partial_{d-2}H_{n,i}$ by $O(n^{5{\zeta}(d-2)})$ balls of radius $n^{4{\zeta}}r_{n}$, denoted $B_{i,\ell}^{(0)}$ say. Replace each ball $B_{i,\ell}^{(0)}$ with a ball $B^{\prime}_{i,\ell}$ with the same centre as $B_{i,\ell}^{(0)}$ and with radius $9dn^{4{\zeta}}r_{n}$. Then cover $B^{\prime}_{i,\ell}$ by $O((n^{4{\zeta}})^{d})$ balls of radius $r_{n}$. Every point in the set on the left side of (3.47) lies in one of these balls of radius $r_{n}$, and the claim follows.
By the definitions of $Q_{n}^{+}$ and $\partial_{d-1}\Gamma_{n}$, and then (3.47), we have
$Q_{n}^{+}\subset\cup_{i=1}^{\sigma^{-}(n)}((\partial_{d-2}H_{n,i})\oplus B(o,8dn^{4{\zeta}}r_{n}))\subset\cup_{i=1}^{\sigma^{-}(n)}\cup_{j=1}^{k_{n,i}}B(x_{i,j},r_{n}),$
so by the union bound,
$\mathbb{P}[F_{n}(Q^{+}_{n},{\cal U}^{\prime}_{n,W_{n}})^{c}]\leq\sum_{i=1}^{\sigma^{-}(n)}\sum_{j=1}^{k_{n,i}}\mathbb{P}[F_{n}(B(x_{i,j},r_{n})\cap A_{n}^{*},{\cal U}^{\prime}_{n,W_{n}})^{c}].$
Thus using Lemma 3.13 and the fact that $\sigma^{-}(n)=O((n^{9{\zeta}}r_{n})^{1-d})$, we obtain that
$\mathbb{P}[F_{n}(Q^{+}_{n},{\cal U}^{\prime}_{n,W_{n}})^{c}]$ $=O((n^{9{\zeta}}r_{n})^{1-d}n^{(9d-10){\zeta}}n^{\varepsilon-1+1/d})$
$=O(n^{-{\zeta}}r_{n}^{1-d}n^{\varepsilon-1+1/d})=O(n^{-{\zeta}/2}).$
Thus $\mathbb{P}[F_{n}(Q^{+}_{n},{\cal U}^{\prime}_{n,W_{n}})^{c}]$ tends to zero. ∎
3.5 Induced coverage process and proof of Proposition 3.10
We shall conclude the proof of Proposition 3.10 by means of a device we refer to as the induced coverage process. This is obtained by taking the parts of $W_{n}$ near the flat parts of $\Gamma_{n}$, along with any Poisson points therein, and rearranging them into a flat region of macroscopic size.
Partition each face $H_{n,i}$, $1\leq i\leq\sigma^{-}(n)$ into a collection of $(d-1)$-dimensional hypercubes of side $n^{3{\zeta}}r_{n}$ contained in $H_{n,i}$ and distant more than $n^{4{\zeta}}r_{n}$ from $\partial_{d-2}H_{n,i}$, together with a ‘border region’ contained within $\partial_{d-2}H_{n,i}\oplus B(o,2n^{4{\zeta}}r_{n})$. Let $P_{n}$ be the union (over all faces) of the boundaries of the $(d-1)$-dimensional hypercubes in this partition (see Figure 4: the $P$ stands for ‘plaid’). Set
$P^{+}_{n}:=[P_{n}\oplus B(o,9n^{{\zeta}}r_{n})]\cap W_{n}^{*}.$ (3.48)
Figure 4: Part of the ‘plaid’ region $P_{n}$ when $d=3$. The outer triangle represents one face $H_{n,i}$, and the part of $P_{n}$ within $H_{n,i}$ is given by the union of the boundaries of the squares. The outer triangle has sides of length $\Theta(n^{9{\zeta}}r_{n})$, while the squares have sides of length $n^{3{\zeta}}r_{n}$. The region between the two triangles has thickness $n^{4{\zeta}}r_{n}$ and is contained in $Q_{n}^{+}$.
Enumerate the $(d-1)$-dimensional hypercubes in the above subdivision of the faces $H_{n,i},1\leq i\leq\sigma^{-}(n)$, as $I^{+}_{n,1},\ldots,I^{+}_{n,\lambda(n)}$. For $1\leq i\leq\lambda(n)$ let $I_{n,i}:=I^{+}_{n,i}\setminus(P_{n}\oplus B(o,n^{\zeta}r_{n}))$, which is a $(d-1)$-dimensional hypercube of side length $(n^{3{\zeta}}-2n^{\zeta})r_{n}$ with the same centre and orientation as $I_{n,i}^{+}$. Writing ||$\cdot$|| ⋅ | below for $(d-1)$-dimensional Hausdorff measure, we claim that the total $(d-1)$-dimensional Hausdorff measure of these $(d-1)$-dimensional hypercubes satisfies
$\lim_{n\to\infty}(|\cup_{i=1}^{\lambda(n)}I_{n,i}|)=|\Gamma|.$ (3.49)
Indeed, for $1\leq i\leq\lambda(n)$ we have $|I_{n,i}|/|I_{n,i}^{+}|=((n^{3{\zeta}}-n^{\zeta})/n^{3{\zeta}})^{d-1}$, which tends to one, so the proportionate amount removed near the boundaries of the $(d-1)$-dimensional hypercubes $I^{+}_{n,i}$ to give $I_{n,i}$ vanishes. Also the ‘border region’ of a face $H_{n,i}$ that is not contained in any of the of the $I^{+}_{n,j}$s has $(d-1)$-dimensional Hausdorff measure that is $O((n^{9{\zeta}}r_{n})^{d-2}n^{4{\zeta}}r_{n})$, so that the total $(d-1)$-dimensional measure of the removed regions near the boundaries of the faces is $O((n^{9{\zeta}}r_{n})^{1-d}\times(n^{9{\zeta}}r_{n})^{d-2}n^{4{\zeta}}r_{n})=O(n^{-5{\zeta}})$, which tends to zero. Thus the claim (3.49) is justified.
For $1\leq i\leq\lambda(n)$, let $S_{n,i}^{+}$, repectively $S_{n,i}$, be a $d$-dimensional cube of side $n^{3{\zeta}}r_{n}$, respectively of side $(n^{3{\zeta}}-2n^{\zeta})r_{n}$, having $I^{+}_{n,i}$, respectively $I_{n,i}$, as its ‘upper face’, i.e. having $I_{n,i}^{+}$ (resp. $I_{n,i}$) as a face and lying within $W_{n}$. We shall verify in Lemma 3.16 that $S_{n,1}^{+},\ldots,S_{n,\lambda(n)}^{+}$ are disjoint.
Define a region $D_{n}\subset\mathbb{R}^{d-1}$ that is approximately a rectilinear hypercube with lower left corner at the origin, and obtained as the union of $\lambda(n)$ disjoint $(d-1)$-dimensional hypercubes of side $n^{3{\zeta}}r_{n}$. We can and do arrange that $D_{n}\subset[0,|\Gamma_{n}|^{1/(d-1)}+n^{3{\zeta}}r_{n}]^{d-1}$ for each $n$, and $|D_{n}|\to|\Gamma|$ as $t\to\infty$. Define the flat slabs (or if $d=2$, flat strips)
$S_{n}:=D_{n}\times[0,(n^{3{\zeta}}-2n^{\zeta})r_{n}];~{}~{}~{}~{}~{}~{}S^{+}_{n}:=D_{n}\times[0,n^{3{\zeta}}r_{n}],$ (3.50)
and denote the lower boundary of $S_{n}$ (that is, the set $D_{n}\times\{0\}$) by $L_{n}$.
Now choose rigid motions $\theta_{n,i}$ of $\mathbb{R}^{d}$, $1\leq i\leq\lambda(n)$, such that under applications of these rigid motions the blocks $S^{+}_{n,i}$ are reassembled to form the slab $S_{n}^{+}$, with the square face $I^{+}_{n,i}$ of the $i$-th block transported to part of the lower boundary $L_{n}$ of $S_{n}$. In other words, choose the rigid motions so that the sets $\theta_{n,i}(S^{+}_{n,i})$, $1\leq i\leq\lambda(n)$, have pairwise disjoint interiors and their union is $S_{n}^{+}$, and also $\theta_{n,i}(I^{+}_{n,i})\subset L_{n}$ for $1\leq i\leq\lambda(n)$.
Recall that $n_{0}$ was chosen shortly before (3.9). Given $n\geq n_{0}$ and $i\leq\lambda(n)$, define $\tilde{\theta}_{n,i}:\mathbb{R}^{d+1}\to\mathbb{R}^{d+1}$ by $\tilde{\theta}_{n,i}(x,r)=(\theta_{n,i}(x),r)$ for all $x\in\mathbb{R}^{d},r\in\mathbb{R}$. By the Restriction, Mapping and Superposition theorems (see e.g. [6]), the point process $\tilde{{\cal U}}_{n,S^{+}_{n}}:=\cup_{i=1}^{\lambda(n)}\tilde{\theta}_{n,i}({\cal U}_{n,W_{n}}\cap(S^{+}_{n,i}\times\mathbb{R}_{+}))$ is a Poisson process in $S_{n}^{+}\times\mathbb{R}_{+}$ with intensity measure $n{\rm Leb}_{d}\otimes\mu_{Y}$ (strictly speaking, with intensity measure $n{\rm Leb}_{d}|_{S_{n}^{+}}\otimes\mu_{Y}$).
We extend $\tilde{\cal U}_{n,S^{+}_{n}}$ to a Poisson process $\tilde{\cal U}_{n,\mathbb{H}}$ in $\mathbb{H}\times\mathbb{R}_{+}$, where $\mathbb{H}:=\mathbb{R}^{d-1}\times[0,\infty)$, as follows. Let $\tilde{\cal U}_{n,\mathbb{H}\setminus S^{+}_{n}}$ be a Poisson process with intensity measure $n{\rm Leb}_{d}\otimes\mu_{Y}$ in $(\mathbb{H}\setminus S^{+}_{n})\times\mathbb{R}_{+}$, independent of ${\cal U}_{n,W_{n}}$, and set
$\tilde{\cal U}_{n,\mathbb{H}}:=\tilde{\cal U}_{n,S_{n}^{+}}\cup\tilde{\cal U}_{n,\mathbb{H}\setminus S_{n}^{+}};~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\tilde{\cal U}^{\prime}_{n,\mathbb{H}}:=\tilde{\cal U}_{n,\mathbb{H}}\cap(\mathbb{R}^{d}\times[0,n^{\zeta}r_{n}]).$ (3.51)
By the Superposition theorem (see e.g. [6, Theorem 3.3]), $\tilde{\cal U}_{n,\mathbb{H}}$ is a homogeneous Poisson process in ×$\mathbb{H}\times[0,\infty)$ with intensity measure $n{\rm Leb}_{d}\times\mu_{Y}$. We call the collection of balls of radius $r_{n}t$ centred on $x$, where $\{(x,t)\}$ are the points of this point process, the induced coverage process.
The next lemma says that the region $P_{n}^{+}$ is covered with high probability. It is needed because locations in $P_{n}^{+}$ lie near the boundary of blocks $S_{n,i}$, so that coverage of these locations by $Z_{n}({\cal U}_{n,A})$ does not necessarily correspond to coverage of their images in the induced coverage process.
Lemma 3.15.
It is the case that $\lim_{t\to\infty}\mathbb{P}[F_{n}(P^{+}_{n},{\cal U}^{\prime}_{n,W_{n}})]=1.$
Proof.
We have $\lambda(n)=O((n^{3{\zeta}}r_{n})^{1-d})$, and for $1\leq i\leq\lambda(n)$, the number of balls of radius $n^{\zeta}r_{n}$ required to cover the boundary of the $(d-1)$-dimensional hypercube $I_{n,i}^{+}$ is $O((n^{3{\zeta}}/n^{\zeta})^{d-2})$. Thus we can take $x_{n,1},\ldots,x_{n,k_{n}}\in\mathbb{R}^{d}$, with $k_{n}=O(r_{n}^{1-d}n^{-{\zeta}-d{\zeta}})$, such that $P_{n}\subset\cup_{i=1}^{k_{n}}B(x_{n,i},n^{\zeta}r_{n})$.
Then $P_{n}^{+}\subset\cup_{i=1}^{k_{n}}B(x_{n,i},10n^{\zeta}r_{n})\cap A_{n}^{*}$, and $B(x_{n,i},10n^{\zeta}r_{n})$ can be covered by $O(n^{d{\zeta}})$ balls of radius $r_{n}$. Hence by taking $\varepsilon={\zeta}/2$ in Lemma 3.13, and using (2.5), we have
$\mathbb{P}[F_{n}(P^{+}_{n},{\cal U}^{\prime}_{n,W_{n}})^{c}]$ $\leq\sum_{i=1}^{k_{n}}\mathbb{P}[F_{n}(B(x_{n,i},10n^{\zeta}r_{n})\cap A_{n}^{*},{\cal U}^{\prime}_{n,W_{n}})^{c}]$
$=O(r_{n}^{1-d}n^{-{\zeta}-d{\zeta}}n^{d{\zeta}}n^{\varepsilon-(d-1)/d})=O(n^{-{\zeta}/2}),$
which tends to zero. ∎
Lemma 3.16.
Suppose $n\geq n_{0}$ and $i<j\leq\lambda(n)$, $i,j\in\mathbb{N}$. Then $(S_{n,i}^{+})^{o}\cap(S_{n,j}^{+})^{o}=\varnothing$.
Proof.
Suppose $(S_{n,i}^{+})^{o}\cap(S_{n,j}^{+})^{o}\neq\varnothing$; we shall obtain a contradiction. Let $x\in(S_{n,i}^{+})^{o}\cap(S_{n,j}^{+})^{o}$. Let $y$ be the closest point in $I^{+}_{n,i}$ to $x$, and $y^{\prime}$ the closest point in $I^{+}_{n,j}$ to $x$. Choose $\ell,m$ such that $I^{+}_{n,i}\subset H_{n,\ell}$ and $I^{+}_{n,j}\subset H_{n,m}$. Then $\ell\neq m$ since if $\ell=m$ we would clearly have $(S_{n,i}^{+})^{o}\cap(S_{n,j}^{+})^{o}=\varnothing$.
Let $J_{n,\ell}\subset\mathbb{R}^{d-1}$ be the image of $H_{n,\ell}$ under projection onto the first $d-1$ coordinates, and write $y=(u,\phi_{n}(u))$ with $u\in J_{n,\ell}$. Let $v\in\partial J_{n,\ell}$, so that $(v,\phi_{n}(v))\in\partial_{d-2}H_{n,\ell}$, By (3.38) and (3.40),
$|\phi_{n}(v)-\phi_{n}(u)|\leq|\phi(v)-\phi(u)|+4Kn^{18{\zeta}}r_{n}^{2}\leq(1/9)\|v-u\|+4Kn^{18{\zeta}}r_{n}^{2}.$
Since $y\in I^{+}_{n,i}$ we have $\|y-(v,\phi_{n}(v))\|\geq\,{\rm dist}(y,\partial H_{n,\ell})\geq n^{4{\zeta}}r_{n}$, so that
$n^{4{\zeta}}r_{n}\leq\|u-v\|+|\phi_{n}(u)-\phi_{n}(v)|\leq(10/9)\|u-v\|+4Kn^{18{\zeta}}r_{n}^{2},$
and hence provided $n$ is large enough, $\|u-v\|\geq n^{4{\zeta}}r_{n}/2$, so that writing $y^{\prime}:=(u^{\prime},\phi_{n}(u^{\prime}))$ we have
$\|y-y^{\prime}\|\geq\|u-u^{\prime}\|\geq\,{\rm dist}(u,\partial J_{n,\ell})\geq n^{4{\zeta}}r_{n}/2.$
But also $\|y-y^{\prime}\|\leq\|y-x\|+\|y^{\prime}-x\|\leq 2n^{3{\zeta}}r_{n}$, and we have our contradiction. ∎
Denote the union of the boundaries (relative to $\mathbb{R}^{d-1}\times\{0\}$) of the lower faces of the blocks making up the strip/slab $S_{n}$, by $C_{n}^{0}$, and the $(9n^{\zeta}r_{n})$-neighbourhood in $\mathbb{H}$ of this region by $C_{n}$ (the $C$ can be viewed as standing for ‘corner region’), i.e.
$C_{n}^{0}:=\cup_{i=1}^{\lambda(n)}\theta_{n,i}(\partial I^{+}_{n,i}),~{}~{}~{}~{}~{}~{}~{}C_{n}:=(C_{n}^{0}\oplus B(o,9n^{\zeta}r_{n}))\cap\mathbb{H}.$ (3.52)
Here $\partial I^{+}_{n,i}$ denotes the relative boundary of $I^{+}_{n,i}$ (relative to the face $H_{n,j}$ containing $I^{+}_{n,i}$).
The next lemma says that the corner region $C_{n}$ is covered with high probability. It is needed because locations in $C_{n}$ lie near the boundaries of the blocks assembled to make the induced coverage process, so that coverage of these locations in the induced coverage process does not necessarily correspond to coverage of their pre-images in the original coverage process.
Lemma 3.17.
It is the case that $\lim_{n\to\infty}\mathbb{P}[F_{n}(C_{n},{\cal U}^{\prime}_{n,\mathbb{H}})]=1$.
Proof.
For each of the $(d-1)$ dimensional hypercubes $\theta_{n,i}(S_{n,i}^{+}),1\leq i\leq\lambda(n)$, the number of balls of radius $n^{\zeta}r_{n}$ required to cover the boundary is $O((n^{3{\zeta}}/n^{\zeta})^{d-2})$. Also $\lambda(n)=O((n^{3{\zeta}}r_{n})^{1-d})$, so we can take points $x_{n,1},\ldots,x_{n,m_{n}}\in L_{n}$, with $m_{n}=O(n^{-{\zeta}-d{\zeta}}r_{n}^{1-d})$, such that $C_{n}^{0}\subset\cup_{i=1}^{m_{n}}B(x_{n,i},n^{\zeta}r_{n})$. Then $C_{n}\subset\cup_{i=1}^{m_{n}}B(x_{n,i},10n^{\zeta}r_{n})$, and $B(x_{i},10n^{\zeta}r_{n})$ can be covered by $O(n^{d{\zeta}})$ balls of radius $r_{n}$. Hence by (3.44) from Lemma 3.13, taking $\varepsilon={\zeta}/2$, we obtain the estimate
$\mathbb{P}[F_{n}(C_{n},{\cal U}^{\prime}_{n,\mathbb{H}})^{c}]=O(n^{d{\zeta}}n^{-{\zeta}-d{\zeta}}r_{n}^{1-d}n^{\varepsilon-(d-1)/d})=O(n^{-{\zeta}/2}),$
which tends to zero. ∎
Lemma 3.18 (Limiting coverage probabilities for the induced coverage process).
With $c_{d,k,Y}$ given at (3.42),
$\lim_{n\to\infty}\mathbb{P}[F_{n}(S_{n},\tilde{\cal U}^{\prime}_{n,\mathbb{H}})]=\lim_{n\to\infty}\mathbb{P}[F_{n}(L_{n},\tilde{\cal U}^{\prime}_{n,\mathbb{H}})]=\exp(-c_{d,k,Y}|\Gamma|e^{-\beta/2}).$ (3.53)
Proof.
The second equality of (3.53) is easily obtained using (3.17) from Lemma 3.5, together with Lemma 3.3.
Recall that $L_{n}=D_{n}\times\{0\}$. Also $\partial L_{n}\subset C_{n}^{0}$, so that $(\partial D_{n}\oplus B_{(d-1)}(o,n^{\zeta}r_{n}))\times[0,2n^{\zeta}r_{n}]\subset C_{n}$, and therefore by (3.18) from Lemma 3.5,
$\mathbb{P}[(F_{n}(L_{n},\tilde{\cal U}^{\prime}_{n,\mathbb{H}})\setminus F_{n}(S_{n},\tilde{\cal U}^{\prime}_{n,\mathbb{H}}))\cap F_{n}(C_{n},\tilde{\cal U}^{\prime}_{n,\mathbb{H}})]\to 0.$
Therefore using also Lemma 3.17 shows that $\mathbb{P}[F_{n}(L_{n},\tilde{\cal U}^{\prime}_{n,\mathbb{H}})\setminus F_{n}(S_{n},\tilde{\cal U}^{\prime}_{n,\mathbb{H}})]\to 0$, and this gives us the rest of (3.53). ∎
Proof of Proposition 3.10.
We shall approximate the event $F_{n}(W^{*}_{n},{\cal U}^{\prime}_{n,W_{n}})$ by events $F_{n}(L_{n},\tilde{\cal U}^{\prime}_{n,\mathbb{H}})$ and $F_{n}(S_{n},\tilde{\cal U}^{\prime}_{n,\mathbb{H}})$, and apply Lemma 3.18. The sets $Q_{n}^{+}$, $P_{n}^{+}$ and $S_{n}$ and were defined at (3.46), (3.48) and (3.50) respectively.
Suppose $F_{n}(Q^{+}_{n}\cup P_{n}^{+},{\cal U}^{\prime}_{n,W_{n}})\setminus F_{n}(W_{n}^{*},{\cal U}^{\prime}_{n,W_{n}})$ occurs, and choose $x\in W_{n}^{*}\setminus(Q_{n}^{+}\cup P_{n}^{+})\setminus Z_{n}({\cal U}^{\prime}_{n,W_{n}})$. Let $y\in\Gamma_{n}$ with $\|y-x\|=\,{\rm dist}(x,\Gamma_{n})$. Then $\|y-x\|\leq 2n^{\zeta}r_{n}$, and since $x\notin Q_{n}^{+}$, we have $\,{\rm dist}(x,\partial_{d-2}\Gamma_{n})\geq 8dn^{4{\zeta}}r_{n}$, and hence $\,{\rm dist}(y,\partial_{d-2}\Gamma_{n})\geq 8dn^{4{\zeta}}r_{n}-2n^{\zeta}r_{n}\geq 3n^{4{\zeta}}r_{n}$, provided $n$ is large enough. Therefore $y$ lies in the interior of the face $H_{n,i}$ for some $i$ and $x-y$ is perpendicular to $H_{n,i}$ (if $y\neq x$). Also, since $x\notin P_{n}^{+}$, $\,{\rm dist}(x,P_{n})\geq 9n^{{\zeta}}r_{n}$, so $\,{\rm dist}(y,P_{n})\geq 7n^{{\zeta}}r_{n}$. Therefore $y\in I_{n,j}$ for some $j$, and $x$ lies in the block $S_{n,j}$. Hence $B(\theta_{n,j}(x),n^{\zeta}r_{n})\cap\mathbb{H}\subset\theta_{n,j}(S^{+}_{n,j})$, and hence by (3.51),
#($\#(\{(z,s)\in\tilde{\cal U}^{\prime}_{n,\mathbb{H}}:\theta_{n,j}(x)\in B(z,r_{n}s)\})=\#(\{(w,t)\in{\cal U}^{\prime}_{n,W_{n}}:x\in B(w,r_{n}t)\})<k,$
so the event $F_{n}(S_{n},\tilde{\cal U}^{\prime}_{n,\mathbb{H}})$ does not occur. Hence
$F_{n}(S_{n},\tilde{\cal U}^{\prime}_{n,\mathbb{H}})\setminus F_{n}(W_{n}^{*},{\cal U}^{\prime}_{n,W_{n}})\subset F_{n}(Q^{+}_{n}\cup P^{+}_{n},{\cal U}^{\prime}_{n,W_{n}})^{c},$
so by Lemmas 3.14 and 3.15, $\mathbb{P}[F_{n}(S_{n},\tilde{\cal U}^{\prime}_{n,\mathbb{H}})\setminus F_{n}(W^{*}_{n},{\cal U}^{\prime}_{n,W_{n}})]\to 0$, and hence using (3.53) we have
$\liminf_{n\to\infty}\mathbb{P}[F_{n}(W_{n}^{*},{\cal U}^{\prime}_{n,W_{n}})]\geq\exp(-c_{d,k,Y}|\Gamma|e^{-\beta/2}).$ (3.54)
Suppose $F_{n}(W^{*}_{n},{\cal U}^{\prime}_{n,W_{n}})\setminus F_{n}(L_{n},\tilde{\cal U}^{\prime}_{n,\mathbb{H}})$ occurs, and choose $y\in L_{n}\setminus Z_{n}(\tilde{\cal U}^{\prime}_{n,\mathbb{H}})$. Take $i\in\{1,\ldots,\lambda(n)\}$ such that $y\in\theta_{n,i}(I_{n,i}^{+})$. Then $\,{\rm dist}(y,\theta_{n,i}(\partial I_{n,i}))\leq n^{\zeta}r_{n}$, since otherwise $\theta_{n,i}^{-1}(y)$ would be a location in $A^{**}_{n}\setminus Z_{n}({\cal U}^{\prime}_{n,W_{n}})$. Thus $y\in C_{n}$ by (3.52), and therefore using Lemma 3.17 yields that
$\mathbb{P}[F_{n}(W^{*}_{n},{\cal U}^{\prime}_{n,W_{n}})\setminus F_{n}(L_{n},\tilde{\cal U}^{\prime}_{n,\mathbb{H}})]\leq\mathbb{P}[F_{n}(C_{n},{\cal U}^{\prime}_{n,W_{n}})^{c}]\to 0.$
Combining this with (3.53) and (3.54) completes the proof. ∎
3.6 Proof of Theorem 2.2
Proposition 3.10 gives the limiting probability of coverage of a polytopal approximation to a region near part of $\partial A$. The next two lemmas show that $\mathbb{P}[F_{n}(W^{*}_{n},{\cal U}^{\prime}_{n,W_{n}})]$ approximates $\mathbb{P}[F_{n}(A^{*}_{n},{\cal U}^{\prime}_{n,A_{n}})]$ (recall the definitions at (3.41)). From this we can deduce that we get the same limiting probability even after dispensing with the polytopal approximation.
Lemma 3.19.
Let $E^{(1)}_{n}:=F_{n}(W_{n}^{*},{\cal U}^{\prime}_{n,W_{n}})\setminus F_{n}(A^{*}_{n},{\cal U}^{\prime}_{n,A_{n}})$. Then $\mathbb{P}[E^{(1)}_{n}]\to 0$ as $n\to\infty$.
Proof.
Let $\varepsilon\in(0,(1/(2d))-9{\zeta})$. Suppose $E^{(1)}_{n}\cap F_{n}(Q_{n}^{+},{\cal U}^{\prime}_{n,A_{n}})$ occurs. Then since ${\cal U}^{\prime}_{n,W_{n}}\subset{\cal U}^{\prime}_{n,A_{n}}$, $\mathbb{R}^{d}\setminus Z_{n}({\cal U}^{\prime}_{n,A_{n}})$ intersects with $A^{*}_{n}\setminus W^{*}_{n}$, and therefore by (3.40), includes locations distant at most $2Kn^{18{\zeta}}r_{n}^{2}$ from $\Gamma_{n}$. Also $\Gamma_{n}\subset Z_{n}({\cal U}^{\prime}_{n,A_{n}})$, since $\Gamma_{n}\subset W^{*}_{n}$.
Pick a location $x\in\overline{A_{n}^{*}\setminus Z_{n}({\cal U}^{\prime}_{n,A_{n}})}$ of minimal distance from $\Gamma_{n}$. Then $x\notin Q_{n}^{+}$, so the nearest point in $\Gamma_{n}$ to $x$ lies in the interior of $H_{n,i}$ for some $i$. We claim that $x$ lies at the intersection of the boundaries of $d$ of the balls making up $Z_{n}({\cal U}^{\prime}_{n,A_{n}})$; this is proved similarly to the similar claim concerning $w$ in the proof of Lemma 3.5. Moreover, $x$ lies in at most $k-1$ of the other balls making up $Z_{n}({\cal U}^{\prime}_{n,A_{n}})$. Also $x$ does not lie in the interior of $W_{n}^{*}$.
Thus if $E^{(1)}_{n}\cap F_{n}(Q_{n}^{+},{\cal U}^{\prime}_{n,A_{n}})$ occurs, there must exist $d$ points $(x_{1},s_{1}),(x_{2},s_{2}),\ldots,$ $(x_{d},s_{d})$ of ${\cal U}_{n,A_{n}}$ such that $\cap_{i=1}^{d}\partial B(x_{i},r_{n}s_{i})$ includes a point in $A_{n}^{*}$ but outside the interior of $W_{n}^{*}$, within distance $2Kn^{18{\zeta}}r_{n}^{2}$ of $\Gamma_{n}$ and in $B(x,r_{n}s)$ for at most $k-1$ of the other points $(x,s)$ of ${\cal U}_{n,A_{n}}$. Hence by Markov’s inequality and the Mecke formula, we obtain that
$\mathbb{P}[E^{(1)}_{n}\cap F_{n}(Q_{n}^{+},{\cal U}^{\prime}_{n,A_{n}})]\leq I_{n,1}+I_{n,2}$ (3.55)
where, taking $Y_{1},\ldots,Y_{d}$ to be independent random variables with the distribution of $Y$, and writing $f_{n,A}(x)$ for $\mathbb{P}[x\in Z_{n}({\cal U}^{\prime}_{n,A})]$ for all $x\in\mathbb{R}^{d}$, and recalling the definition of $q_{n}(\cdot)$ at (3.15), we set
$I_{n,1}:=$ $n^{d}\int_{\mathbb{R}^{d}}\cdots\int_{\mathbb{R}^{d}}\mathbb{E}\,[h_{n}((x_{1},Y_{1}),(x_{1}+y_{2},Y_{2}),\ldots,(x_{1}+y_{d},Y_{d}))$
×$\times{\bf 1}\{q_{n}((x_{1},Y_{1}),(x_{1}+y_{2},Y_{2}),\ldots,(x_{1}+y_{d},Y_{d}))\in A_{n}^{*}\cap(\Gamma_{n}\oplus B(o,2Kn^{18{\zeta}}r_{n}^{2}))\}$
×$\times(1-f_{n,A}(q_{n}((x_{1},Y_{1}),(x_{1}+y_{2},Y_{2}),\ldots,(x_{1}+y_{d},Y_{d}))))]dy_{d}\cdots dy_{2}dx_{1},$
and $I_{n,2}$ is defined similarly with $p_{n}(\cdot)$ replacing $q_{n}(\cdot)$. Then changing variables $y_{i}\mapsto r_{n}^{-1}y_{i}$ we have
$I_{n,1}=$ $n^{d}r_{n}^{d(d-1)}\mathbb{E}\,\Big{[}\int_{\mathbb{R}^{d}}\cdots\int_{\mathbb{R}^{d}}h_{n}((o,Y_{1}),(r_{n}y_{2},Y_{2}),\ldots,(r_{n}y_{d},Y_{d}))$
×$\times{\bf 1}\{x_{1}+q_{n}((o,Y_{1}),(r_{n}y_{2},Y_{2}),\ldots,(r_{n}y_{d}Y_{d}))\in A_{n}^{*}\cap(\Gamma_{n}\oplus B(o,2Kn^{18{\zeta}}r_{n}^{2}))\}$
×$\times(1-f_{n,A}(x_{1}+q_{n}((o,Y_{1}),(r_{n}y_{2},Y_{2}),\ldots,(r_{n}y_{d},Y_{d}))))dy_{d}\cdots dy_{2}dx_{1}\Big{]}.$
By Lemma 3.13, the last factor of $1-f_{n,A}(\cdot)$ is $O(n^{\varepsilon-1+1/d})$ whenever the indicator function in the previous factor is 1. Also using Fubini’s theorem we can take the integral over $x_{1}$ inside all the other integrals and integrate it out, getting a factor of $O(n^{18{\zeta}}r_{n}^{2})$. Thus we obtain for a suitable constant $c$ that
$I_{n,1}\leq cn^{d}r_{n}^{d(d-1)}n^{18{\zeta}+\varepsilon-1+1/d}r_{n}^{2}\mathbb{E}\,\Big{[}\int_{\mathbb{R}^{d}}\cdots\int_{\mathbb{R}^{d}}h((o,r_{n}Y_{1}),(r_{n}y_{2},r_{n}Y_{2}),\ldots,(r_{n}y_{d},r_{n}Y_{d}))$
$dy_{d}\cdots dy_{2}\Big{]}$
[…]
In the last line the expectation is finite by Lemma 3.4 and our moment condition on $Y$. Also $(nr_{n}^{d})^{d}=O(n^{\varepsilon})$ by (2.5). Thus $I_{n,1}=O(n^{2\varepsilon+18{\zeta}-1/d})$ so $I_{n,1}\to 0$ as $n\to\infty$, and by an identical argument the same holds for $I_{n,2}$. Also $\mathbb{P}[F_{n}(Q_{n}^{+},{\cal U}^{\prime}_{n,A_{n}})]\to 1$ by Lemma 3.14, so using (3.55) we obtain that $\mathbb{P}[E^{(1)}_{n}]\to 0$, as required. ∎
Lemma 3.20.
Let $E^{(2)}_{n}:=F_{n}(A_{n}^{*},{\cal U}^{\prime}_{n,A_{n}})\setminus F_{n}(W_{n}^{*},{\cal U}^{\prime}_{n,W_{n}})$. Then $\lim_{n\to\infty}\mathbb{P}[E^{(2)}_{n}]=0$.
Proof.
If the event $E^{(2)}_{n}$ occurs, then since $W^{*}_{n}\subset A_{n}^{*}$, the set $W^{*}_{n}\cap Z_{n}({\cal U}^{\prime}_{n,A_{n}})\setminus Z_{n}({\cal U}^{\prime}_{n,W_{n}})$ is nonempty. Hence there exists $(x,s)\in{\cal U}^{\prime}_{n,A_{n}}\setminus{\cal U}^{\prime}_{n,W_{n}}$ with $B(x,r_{n}s)\cap W^{*}_{n}\setminus Z_{n}({\cal U}^{\prime}_{n,W_{n}})\neq\varnothing$. Therefore
$E^{(2)}_{n}\subset F_{n}(Z_{n}({\cal U}^{\prime}_{n,A_{n}}\setminus{\cal U}^{\prime}_{n,W_{n}})\cap W_{n}^{*},{\cal U}^{\prime}_{n,W_{n}})^{c}.$ (3.56)
Let $\varepsilon\in(0,(1/(2d))-(9+d/2){\zeta})$. Let ${\cal Q}_{n}:={\cal U}^{\prime}_{n,A_{n}}\setminus{\cal U}^{\prime}_{n,W_{n}}$. Then ${\cal U}^{\prime}_{n,W_{n}}$ and ${\cal Q}_{n}$ are independent Poisson processes with intensity measures $n{\rm Leb}_{d}\otimes\mu_{Y}$ in $W_{n}\times[0,n^{\zeta}]$, $(A_{n}\setminus W_{n})\times[0,n^{\zeta}]$ respectively. By Lemma 3.13 and the union bound, there is a constant $c$ such that for any $m\in\mathbb{N}$ and any set of $m$ points $(x_{1},t_{1}),\ldots,(x_{m},t_{m})$ in $\mathbb{R}^{d}\times[0,n^{\zeta}]$, we have
$\mathbb{P}\left[F_{n}(\cup_{i=1}^{m}B(x_{i},r_{n}t_{i})\cap W_{n}^{*},{\cal U}^{\prime}_{n,W_{n}})^{c}\right]\leq cmn^{d{\zeta}}n^{\varepsilon-(d-1)/d}.$
Let $N_{n}:={\cal Q}_{n}(\mathbb{R}^{d}\times[0,n^{\zeta}])$. By Lemma 3.11(a), $\mathbb{E}\,[N_{n}]=O(n^{1+18{\zeta}}r_{n}^{2})$, so that by conditioning on ${\cal Q}_{n}$ we have
$\mathbb{P}[F_{n}(\cup_{(x,t)\in{\cal Q}_{n}}B(x,r_{n}t)\cap W_{n}^{*},{\cal U}^{\prime}_{n,W_{n}})^{c}]$ $\leq cn^{\varepsilon+d{\zeta}-1+1/d}\mathbb{E}\,[N_{n}]$
$=O(n^{(18+d){\zeta}+2\varepsilon-1/d}),$
which tends to zero by the choice of $\varepsilon$. Hence by (3.56), $\mathbb{P}[E^{(2)}_{n}]\to 0$. ∎
To complete the proof of Theorem 2.10, we shall break $\partial A$ into finitely many pieces, with each piece contained in a single chart. We would like to write the probability that all of $\partial A$ is covered as the product of probabilities for each piece, but to achieve the independence needed for this, we need to remove a region near the boundary of each piece. By separate estimates we can show the removed regions are covered with high probability, and this is the content of the next lemma.
With $\Gamma$ and $\partial\Gamma$ as in Section 3.4, define the sets ${\Delta}_{n}:=\partial\Gamma\oplus B(o,n^{29{\zeta}}r_{n})$ and ${\Delta}_{n}^{+}:=\partial\Gamma\oplus B(o,n^{49{\zeta}}r_{n}).$
Lemma 3.21.
It is the case that $\lim_{n\to\infty}F_{n}({\Delta}_{n}^{+}\cap A,{\cal U}^{\prime}_{n,A})=1$.
Proof.
Let $\varepsilon\in(0,(1/d)-98{\zeta})$. Since we assume $\kappa(\partial\Gamma,r)=O(r^{2-d})$ as $r\downarrow 0$, for each $n$ we can take $x_{n,1},\ldots,x_{n,k(n)}\in\mathbb{R}^{d}$ with $\partial\Gamma\subset\cup_{i=1}^{k(n)}B(x_{n,i},n^{49{\zeta}}r_{n})$, and with $k(n)=O((n^{49{\zeta}}r_{n})^{2-d})$. Then ${\Delta}_{n}^{+}\subset\cup_{i=1}^{k(n)}B(x_{n,i},2n^{49{\zeta}}r_{n})$. For each $i\in\{1,\ldots,k(n)\}$, we can cover the ball $B(x_{n,i},2n^{49{\zeta}}r_{n})$ with $O(n^{49d{\zeta}})$ smaller balls of radius $r_{n}$. Then we end up with balls of radius $r_{n}$, denoted $B_{n,1},\ldots,B_{n,m(n)}$ say, such that ${\Delta}_{n}^{+}\subset\cup_{i=1}^{m(n)}B_{n,i}$ and $m(n)=O(r_{n}^{2-d}n^{49{\zeta}(3-d)})=O(r_{n}^{2-d}n^{98{\zeta}})$. By (3.45) from Lemma 3.13, and the union bound,
$\mathbb{P}[\cup_{i=1}^{m(n)}(F_{n}(B_{n,i}\cap A,{\cal U}^{\prime}_{n,A})^{c})]=O(r_{n}^{2-d}n^{98{\zeta}+\varepsilon-1+1/d})=O(n^{98{\zeta}+\varepsilon-1/d}),$
which tends to zero. ∎
Given $n>0$, define the sets $\Gamma^{(n^{29{\zeta}}r_{n})}:=\Gamma\setminus{\Delta}_{n}$ and
$\Gamma^{(n^{29{\zeta}}r_{n})}_{n^{\zeta}r_{n}}:=(\Gamma^{(n^{29{\zeta}}r_{n})}\oplus B(o,n^{\zeta}r_{n}))\cap A;~{}~{}~{}~{}~{}\Gamma_{n^{\zeta}r_{n}}:=(\Gamma\oplus B(o,n^{\zeta}r_{n}))\cap A,$
and define the event $F_{n}^{\Gamma}:=F_{n}(\Gamma^{(n^{29{\zeta}}r_{n})}_{n^{\zeta}r_{n}},{\cal U}^{\prime}_{n,A})$.
Note that the definition of $F_{n}^{\Gamma}$ does not depend on the choice of chart. This fact will be needed for the last stage of the proof of Theorem 2.10. Lemma 3.24 below shows that $\mathbb{P}[F_{n}^{\Gamma}]$ is well approximated by $\mathbb{P}[F_{n}(A_{n}^{*},{\cal U}^{\prime}_{n,A_{n}})]$ and we have already determined the limiting behaviour of the latter. We prepare for the proof of Lemma 3.24 with two geometrical lemmas.
Lemma 3.22.
For all large enough $n$, it is the case that $\Gamma_{n^{\zeta}r_{n}}^{(n^{29{\zeta}}r_{n})}\subset A_{n}^{*}$.
Proof.
Let $x\in\Gamma^{(n^{29{\zeta}}r_{n})}_{n^{\zeta}r_{n}}$, and take $y\in\Gamma^{(n^{29{\zeta}}r_{n})}$ with $\|x-y\|\leq n^{\zeta}r_{n}$. Writing $y=(u,\phi(u))$ with $u\in U$, we claim that $\,{\rm dist}(u,\partial U)\geq(1/2)n^{29{\zeta}}r_{n}$. Indeed, if we had $\,{\rm dist}(u,\partial U)<(1/2)n^{29{\zeta}}r_{n}$, then we could take $w\in\partial U$ with $\|u-w\|<(1/2)n^{29{\zeta}}r_{n}$. Then $(w,\phi(w))\in\partial\Gamma$ and by (3.38), $|\phi(w)-\phi(u)|\leq(1/4)n^{29{\zeta}}r_{n}$, so
$\|(u,\phi(u))-(w,\phi(w))\|\leq\|u-w\|+|\phi(u)-\phi(w)|\leq(3/4)n^{29{\zeta}}r_{n},$
contradicting the assumption that $y\in\Gamma^{(n^{29{\zeta}}r_{n})}$, so the claim is justified.
Writing $x=(v,s)$ with $v\in\mathbb{R}^{d-1}$, and $s\in\mathbb{R}$, we have $\|v-u\|\leq\|x-y\|\leq n^{\zeta}r_{n}$, so $\,{\rm dist}(v,\partial U)\geq(1/2)n^{29{\zeta}}r_{n}-n^{\zeta}r_{n}$, and hence $v\in U_{n}^{-}$, provided $n$ is big enough ($U_{n}^{-}$ was defined shortly after (3.39).) Also $|\phi(v)-\phi(u)|\leq n^{\zeta}r_{n}/4$ by (3.38), so $|\phi_{n}(v)-\phi(u)|\leq n^{\zeta}r_{n}/2$, provided $n$ is big enough, by (3.40). Also $|s-\phi(u)|\leq\|x-y\|\leq n^{\zeta}r_{n}$, so $|s-\phi_{n}(v)|\leq(3/2)n^{\zeta}r_{n}.$ Therefore $x\in A^{*}_{n}$ by (3.41). ∎
Lemma 3.23.
For all large enough $n$, we have (a) $[A_{n}^{*}\oplus B(o,4n^{\zeta}r_{n})]\cap A\subset A_{n}$, and (b) $[A_{n}^{*}\oplus B(o,4n^{\zeta}r_{n})]\cap\partial A\subset\Gamma$, and (c) $[\Gamma_{n^{\zeta}r_{n}}^{(n^{29{\zeta}}r_{n})}\oplus B(o,4n^{\zeta}r_{n})]\cap\partial A\subset\Gamma$.
Proof.
Let $x\in A_{n}^{*}$. Write $x=(u,z)$ with $u\in U_{n}^{-}$ and $\phi_{n}(u)-3n^{\zeta}r_{n}/2\leq z\leq\phi(u)$.
Let $y\in B(x,4n^{\zeta}r_{n})\cap A$, and write $y=(v,s)$ with $v\in\mathbb{R}^{d-1}$ and $s\in\mathbb{R}$. Then $\|v-u\|\leq 4n^{\zeta}r_{n}$ so provided $n$ is big enough, $v\in U_{n}$. Also $|s-z|\leq 4n^{\zeta}r_{n}$, and $|\phi(v)-\phi(u)|\leq n^{\zeta}r_{n}$ by (3.38), so
$|s-\phi(v)|\leq|s-z|+|z-\phi(u)|+|\phi(u)-\phi(v)|\leq 4n^{\zeta}r_{n}+2n^{\zeta}r_{n}+n^{\zeta}r_{n},$
and since $y\in A$, by (3.36) and (3.37) we must have $0\leq s\leq\phi(v)$, provided $n$ is big enough. Therefore $y=(v,s)\in A_{n}$, which gives us (a).
If also $y\in\partial A$, then $\phi(v)=s$, so $y\in\Gamma$. Hence we have part (b). Then by Lemma 3.22 we also have part (c). ∎
Lemma 3.24.
It is the case that $\mathbb{P}[F_{n}^{\Gamma}\triangle F_{n}(A_{n}^{*},{\cal U}^{\prime}_{n,A_{n}})]\to 0$ as $n\to\infty$.
Proof.
Since $\Gamma_{n^{\zeta}r_{n}}^{(n^{29{\zeta}}r_{n})}\subset A_{n}^{*}$ by Lemma 3.22, and moreover ${\cal U}^{\prime}_{n,A_{n}}\subset{\cal U}^{\prime}_{n,A}$, it follows that $F_{n}(A_{n}^{*},{\cal U}^{\prime}_{n,A_{n}})\subset F_{n}(\Gamma^{(n^{29{\zeta}}r_{n})}_{n^{\zeta}r_{n}},{\cal U}^{\prime}_{n,A})=F_{n}^{\Gamma}$. Therefore it suffices to prove that
$\mathbb{P}[F_{n}(\Gamma^{(n^{29\xi}r_{n})}_{n^{\zeta}r_{n}},{\cal U}^{\prime}_{n,A})\setminus F_{n}(A_{n}^{*},{\cal U}^{\prime}_{n,A_{n}})]\to 0.$ (3.57)
Let $\varepsilon>0$. Suppose event $F_{n}(\Gamma^{(n^{29\xi}r_{n})}_{n^{\zeta}r_{n}},{\cal U}^{\prime}_{n,A})\cap F_{n}({\Delta}_{n}^{+}\cap A,{\cal U}^{\prime}_{n,A})\setminus F_{n}(A_{n}^{*},{\cal U}^{\prime}_{n,A_{n}})$ occurs. Choose $x\in A_{n}^{*}\setminus Z_{n}({\cal U}^{\prime}_{n,A_{n}})$. Then by Lemma 3.23(a), $B(x,n^{\zeta}r_{n})\cap A\subset A_{n}$. Hence ${\cal U}^{\prime}_{n,A}\cap(B(x,n^{\zeta}r_{n})\times\mathbb{R}_{+})\subset{\cal U}^{\prime}_{n,A_{n}}$, and therefore $x\notin Z_{n}({\cal U}^{\prime}_{n,A})$.
Since we are assuming $F_{n}(\Gamma^{(n^{29{\zeta}}r_{n})}_{n^{\zeta}r_{n}},{\cal U}^{\prime}_{n,A})$ occurs, we therefore have $\,{\rm dist}(x,\Gamma^{(n^{29{\zeta}}r_{n})})>n^{\zeta}r_{n}$. Since we also assume $F_{n}({\Delta}_{n}^{+}\cap A,{\cal U}^{\prime}_{n,A})$, we also have $\,{\rm dist}(x,\partial\Gamma)\geq n^{49{\zeta}}r_{n}$ and therefore $\,{\rm dist}(x,{\Delta}_{n})=\,{\rm dist}(x,\partial\Gamma\oplus B(o,n^{29{\zeta}}r_{n}))\geq n^{29{\zeta}}r_{n}$. Hence
$\,{\rm dist}(x,\Gamma)\geq\min(\,{\rm dist}(x,\Gamma^{(n^{29{\zeta}}r_{n})}),\,{\rm dist}(x,\partial\Gamma\oplus B(o,n^{29{\zeta}}r_{n}))>n^{\zeta}r_{n}.$
Moreover, by Lemma 3.23(b), $\,{\rm dist}(x,(\partial A)\setminus\Gamma)>n^{\zeta}r_{n}$. Thus $\,{\rm dist}(x,\partial A)>n^{\zeta}r_{n}$. Moreover, $\,{\rm dist}(x,\partial A)\leq\,{\rm dist}(x,\Gamma)\leq 2n^{\zeta}r_{n}$ because $x\in A_{n}^{*}$, and therefore $x\notin A^{[\varepsilon]}$ (provided $n$ is large enough) since $\overline{A^{[\varepsilon]}}$ is compact and contained in $A^{o}$ (the set $A^{[\varepsilon]}$ was defined in Section 2.) Therefore the event $F_{n}(A^{(n^{\zeta}r_{n})}\setminus A^{[\varepsilon]},{\cal U}^{\prime}_{n,A})^{c}$ occurs. Thus, for large enough $n$ we have the event inclusion
$F_{n}(\Gamma^{(n^{29{\zeta}}r_{n})}_{r_{n}},{\cal U}^{\prime}_{n,A})\cap F_{n}({\Delta}_{n}^{+}\cap A,{\cal U}^{\prime}_{n,A})\setminus F_{n}(A_{n}^{*},{\cal U}^{\prime}_{n,A_{n}})\subset F_{n}(A^{(n^{\zeta}r_{n})}\setminus A^{[\varepsilon]},{\cal U}^{\prime}_{n,A})^{c}.$ (3.58)
By (2.5),
$\lim_{n\to\infty}(\omega_{d}\mathbb{E}\,[Y^{d}]nr_{n}^{d}-\log n-(d+k-2)\log\log n)=\begin{cases}\beta&{\rm if}~{}d=2,k=1\\ +\infty&{\rm otherwise.}\end{cases}$ (3.59)
Hence by Lemma 3.2 applied to the set $A\setminus A^{[\varepsilon]}$, and Lemma 3.3 we have that
$\liminf_{n\to\infty}\mathbb{P}[F_{n}(A^{(n^{\zeta}r_{n})}\setminus A^{[\varepsilon]},{\cal U}^{\prime}_{n,A})]=\liminf_{n\to\infty}\mathbb{P}[F_{n}(A^{(n^{\zeta}r_{n})}\setminus A^{[\varepsilon]},{\cal U}^{\prime}_{n,\mathbb{R}^{d}})]$
$\geq\liminf_{n\to\infty}\mathbb{P}[F_{n}(A\setminus A^{[\varepsilon]},{\cal U}_{n,\mathbb{R}^{d}})]$
=$=\begin{cases}\exp\Big{(}-c_{d}\Big{(}\frac{(\mathbb{E}\,[Y^{d-1}])^{d}}{(\mathbb{E}\,[Y^{d}])^{d-1}}\Big{)}|A\setminus A^{[\varepsilon]}|e^{-\beta}\Big{)}&{\rm if~{}}d=2,k=1\\ 1&{\rm otherwise.}\end{cases}$ (3.60)
Therefore since $\varepsilon$ can be arbitrarily small and $|A\setminus A^{[\varepsilon]}|\to 0$ as $\varepsilon\downarrow 0$, the event displayed on the left hand side of (3.58) has probability tending to zero. Then using Lemma 3.21, we have (3.57), which completes the proof. ∎
Corollary 3.25.
It is the case that $\lim_{n\to\infty}\mathbb{P}[F_{n}^{\Gamma}]=\exp(-c_{d,k,Y}|\Gamma|e^{-\beta/2}).$
Proof.
By Lemmas 3.19 and 3.20, $\mathbb{P}[F_{n}(W^{*}_{n},{\cal U}^{\prime}_{n,W_{n}})\triangle F_{n}(A_{n}^{*},{\cal U}^{\prime}_{n,A_{n}})]\to 0$. Then by Lemma 3.24, $\mathbb{P}[F_{n}^{\Gamma}\triangle F_{n}(W^{*}_{n},{\cal U}^{\prime}_{n,W_{n}})]\to 0$, and now the result follows by Proposition 3.10. ∎
Proof of Theorem 2.2.
Let $x_{1},\ldots,x_{J}$ and $r(x_{1}),\ldots,r(x_{J})$ be as described at (3.35). Set $\Gamma_{1}:=B(x_{1},r(x_{1}))\cap\partial A$, and for $j=2,\ldots,J$, let
$\Gamma_{j}:=\overline{B(x_{j},r(x_{j}))\cap\partial A\setminus\cup_{i=1}^{j-1}B(x_{i},r(x_{i}))},$
and $\partial\Gamma_{i}:=\Gamma_{i}\cap\overline{\partial A\setminus\Gamma_{i}}$. Then $\Gamma_{1},\ldots,\Gamma_{J}$ comprise a finite collection of closed sets in $\partial A$ with disjoint interiors, each of which satisfies $\kappa(\partial\Gamma_{i},r)=O(r^{2-d})$ as $r\downarrow 0$, and is contained in a single chart $B(x_{j},r(x_{j}))$, and with union $\partial A$. For $1\leq i\leq J$, define $F_{n}^{\Gamma_{i}}$ analogously to $F_{n}^{\Gamma}$, that is, $F_{n}^{\Gamma_{i}}:=F_{n}(\Gamma_{i,n^{{\zeta}}r_{n}}^{(n^{29{\zeta}}r_{n})},{\cal U}^{\prime}_{n,A})$ with
$\Gamma_{i,n^{\zeta}r_{n}}^{(n^{29{\zeta}}r_{n})}:=\left(\left[\Gamma_{i}\setminus((\partial\Gamma_{i})\oplus B(o,n^{29{\zeta}}r_{n}))\right]\oplus B(o,n^{\zeta}r_{n})\right)\cap A.$
First we claim that the following event inclusion holds:
$\cap_{i=1}^{J}F_{n}^{\Gamma_{i}}\cap F_{n}(A^{(n^{\zeta}r_{n})},{\cal U}^{\prime}_{n,A})\setminus F_{n}(A,{\cal U}^{\prime}_{n,A})\subset\left(\cap_{i=1}^{J}F_{n}([(\partial\Gamma_{i})\oplus B(o,n^{49{\zeta}}r_{n})]\cap A,{\cal U}^{\prime}_{n,A})\right)^{c}.$
Indeed, suppose $\cap_{i=1}^{J}F_{n}^{\Gamma_{i}}\cap F_{n}(A^{(n^{\zeta}r_{n})},{\cal U}^{\prime}_{n,A})\setminus F_{n}(A,{\cal U}^{\prime}_{n,A})$ occurs, and choose $x\in A\setminus Z_{n}({\cal U}^{\prime}_{n,A})$. Then $\,{\rm dist}(x,\partial A)\leq n^{\zeta}r_{n}$ since we assume $F_{n}(A^{(n^{\zeta}r_{n})},{\cal U}^{\prime}_{n,A})$ occurs. Then for some $i\in\{1,\ldots,J\}$ and some $y\in\Gamma_{i}$ we have $\|x-y\|\leq n^{\zeta}r_{n}$. Since we assume $F_{n}^{\Gamma_{i}}$ occurs, we have $x\notin\Gamma_{i,n^{\zeta}r_{n}}^{(n^{29{\zeta}}r_{n})}$, and hence $\,{\rm dist}(y,\partial\Gamma_{i})\leq n^{29{\zeta}}r_{n}$, so $\,{\rm dist}(x,\partial\Gamma_{i})<n^{49{\zeta}}r_{n}$. Therefore $F_{n}([(\partial\Gamma_{i})\oplus B(o,n^{49{\zeta}}r_{n})]\cap A,{\cal U}^{\prime}_{n,A})$ fails to occur, justifying the claim.
By the preceding claim and and the union bound,
$\mathbb{P}[F_{n}(A,{\cal U}^{\prime}_{n,A})]$ $\leq\mathbb{P}[\cap_{i=1}^{J}F_{n}^{\Gamma_{i}}\cap F_{n}(A^{(n^{\zeta}r_{n})},{\cal U}^{\prime}_{n,A})]$
$\leq\mathbb{P}[F_{n}(A,{\cal U}^{\prime}_{n,A})]+\sum_{i=1}^{J}\mathbb{P}[F_{n}([(\partial\Gamma_{i})\oplus B(o,n^{49{\zeta}}r_{n})]\cap A,{\cal U}^{\prime}_{n,A})^{c}].$
By Lemma 3.21, $\mathbb{P}[F_{n}([(\partial\Gamma_{i})\oplus B(o,n^{49{\zeta}}r_{n})]\cap A,{\cal U}^{\prime}_{n,A})]\to 1$ for each $i$. Therefore
$\lim_{n\to\infty}\mathbb{P}[F_{n}(A,{\cal U}^{\prime}_{n,A})]=\lim_{n\to\infty}\mathbb{P}[\cap_{i=1}^{J}F_{n}^{\Gamma_{i}}\cap F_{n}(A^{(n^{\zeta}r_{n})},{\cal U}^{\prime}_{n,A})],$ (3.61)
provided the last limit exists. By Corollary 3.25, we have for each $i$ that
$\lim_{t\to\infty}(\mathbb{P}[F^{\Gamma_{i}}_{n}])=\exp(-c_{d,k,Y}|\Gamma_{i}|e^{-\beta/2}).$ (3.62)
Also, we claim that for large enough $n$ the events $F_{n}^{\Gamma_{1}}$, …, $F_{n}^{\Gamma_{J}}$ are mutually independent. Indeed, given distinct $i,j\in\{1,\ldots,J\}$, if $x\in\Gamma_{i,n^{\zeta}r_{n}}^{(n^{29{\zeta}}r_{n})}$ and $y\in\Gamma_{j,n^{\zeta}r_{n}}^{(n^{29{\zeta}}r_{n})}$, then we can take $y^{\prime}\in\Gamma_{j}\setminus(\partial\Gamma_{j}\oplus B(o,n^{29{\zeta}}r_{n}))$ with $\|y^{\prime}-y\|\leq n^{\zeta}r_{n}$. If $\|x-y\|\leq 3n^{\zeta}r_{n}$ then by the triangle inequality $\|x-y^{\prime}\|\leq 4n^{\zeta}r_{n}$, but since $y^{\prime}\notin\Gamma_{i}$, this would contradict Lemma 3.23(c). Therefore $\|x-y\|>3n^{\zeta}r_{n}$, and hence the $n^{\zeta}r_{n}$-neighbourhoods of $\Gamma_{i,n^{\zeta}r_{n}}^{(n^{29{\zeta}}r_{n})}$ and of $\Gamma_{j,n^{\zeta}r_{n}}^{(n^{29{\zeta}}r_{n})}$ are disjoint. This gives us the independence claimed.
Now observe that $F_{n}(A^{(n^{\zeta}r_{n})},{\cal U}^{\prime}_{n,A})\subset F_{n}(A^{(4n^{\zeta}r_{n})},{\cal U}^{\prime}_{n,A})$. We claim that
$\mathbb{P}[F_{n}(A^{(4n^{\zeta}r_{n})},{\cal U}^{\prime}_{n,A})\setminus F_{n}(A^{(n^{\zeta}r_{n})},{\cal U}^{\prime}_{n,A})]\to 0~{}~{}~{}{\rm as}~{}n\to\infty.$ (3.63)
Indeed, given $\varepsilon>0$, for large $n$ the probability on the left side of (3.63) is bounded by $\mathbb{P}[F_{n}(A^{(n^{\zeta}r_{n})}\setminus A^{[\varepsilon]},{\cal U}^{\prime}_{n,A})^{c}]$, and by (3.60) the limsup of the latter probability can be made arbitrarily small by the choice of $\varepsilon$. Hence by Lemma 3.2 and (3.59), Lemma 3.3 and the fact that $c_{2}=1$,
$\lim_{n\to\infty}\mathbb{P}[F_{n}(A^{(4n^{\zeta}r_{n})},{\cal U}^{\prime}_{n,A})]$ $=\lim_{n\to\infty}\mathbb{P}[F_{n}(A^{(n^{\zeta}r_{n})},{\cal U}^{\prime}_{n,A})]$
$=\lim_{n\to\infty}\mathbb{P}[F_{n}(A^{(n^{\zeta}r_{n})},{\cal U}^{\prime}_{n,\mathbb{R}^{d}})]$
=$=\begin{cases}\exp\Big{(}-\Big{(}\frac{(\mathbb{E}\,[Y])^{2}}{\mathbb{E}\,[Y^{2}]}\Big{)}|A|e^{-\beta}\Big{)}&{\rm if}~{}d=2,k=1\\ 1&{\rm otherwise}.\end{cases}$ (3.64)
Moreover, by (3.61) and (3.63),
$\lim_{n\to\infty}\mathbb{P}[F_{n}(A,{\cal U}^{\prime}_{n,A})]=\lim_{n\to\infty}\mathbb{P}[\cap_{i=1}^{J}F_{n}^{\Gamma_{i}}\cap F_{n}(A^{(4n^{\zeta}r_{n})},{\cal U}^{\prime}_{n,A})],$ (3.65)
provided the last limit exists. However, the events in the right hand side of (3.65) are mutually independent, so using (3.62), (3.42) and (3.64), we obtain that
$\lim_{n\to\infty}\mathbb{P}[F_{n}(A,{\cal U}^{\prime}_{n,A})]=\exp\Big{(}-c_{d,k}\Big{(}\frac{(\mathbb{E}\,[Y^{d-1}])^{d-1}}{(\mathbb{E}\,[Y^{d}])^{d-2+1/d}}\Big{)}|\partial A|e^{-\beta/2}$
$-\Big{(}\frac{\mathbb{E}\,[Y]^{2}}{\mathbb{E}\,[Y^{2}]}\Big{)}|A|e^{-\beta}{\bf 1}_{\{(d,k)=(2,1)\}}\Big{)}.$
By Lemma 3.3, if we replace ${\cal U}^{\prime}_{n,A}$ with ${\cal U}_{n,A}$ on the left, we get the same limit, i.e. (2.6) if $d\geq 3$ and (2.1) if $d=2$. ∎
4 Proof of results for the J-M model
Lemma 4.1 (Scaling lemma).
Suppose $L>0$ and $\rho=L^{d+1}$. Then $L\tilde{T}_{\rho}$ has the same distribution as $\tilde{\tau}_{L}$. Moreover $LT_{\rho}$ has the same distribution as $\tau_{L}$.
Proof.
For any $t>0$, setting $t^{\prime}=Lt$, we have
$A\subset\cup_{(x,s)\in{\cal H}_{\rho}\cap(\mathbb{R}^{d}\times[0,t])}B(x,t-s)\Leftrightarrow A_{L}\subset\cup_{(y,u)\in L{\cal H}_{\rho}\cap(\mathbb{R}^{d}\times[0,t^{\prime}])}B(y,t^{\prime}-u)$
so that by (2.8),
$L\tilde{T}_{\rho}=\inf\{t^{\prime}:A_{L}\subset\cup_{(y,u)\in L{\cal H}_{\rho}\cap(\mathbb{R}^{d}\times[0,t^{\prime}])}B(y,t^{\prime}-u)\}.$
By our choice of $\rho$ and the Mapping theorem (see e.g. [6]), $L{\cal H}_{\rho}$ is a homogeneous Poisson process of intensity $1$ on $\mathbb{R}^{d}\times\mathbb{R}_{+}$, and therefore by the definition (2.10), $L\tilde{T}_{\rho}$ has the same distribution as $\tilde{\tau}_{L}$.
The proof that $LT_{\rho}$ has the same distribution as $\tau_{L}$ is similar. ∎
Proof of Proposition 2.8.
Suppose for some $\beta\in\mathbb{R}$ that $(t_{\rho})_{\rho>0}$ satisfies
$\lim_{\rho\to\infty}\left(\omega_{d}\rho t_{\rho}^{d+1}-d\log\rho-d^{2}\log\log\rho\right)=\beta,$ (4.1)
i.e. $t_{\rho}=\left(\left(d\log\rho+d^{2}\log\log\rho+\beta+o(1)\right)/(\omega_{d}\rho)\right)^{1/(d+1)}$.
Write ${\cal H}_{\rho}\cap(\mathbb{R}^{d}\times[0,t_{\rho}])=\{(X_{i},S_{i})\}_{i\in\mathbb{N}}$. Then the point process ${\cal P}_{\rho t_{\rho}}:=\{X_{i}\}_{i\in\mathbb{N}}$ is a homogeneous Poisson process of intensity $\rho t_{\rho}$ in $\mathbb{R}^{d}$.
The balls $B(X_{i},t_{\rho}-S_{i})$, $i\in\mathbb{N},$ form a SPBM in $\mathbb{R}^{d}$, and in the notation of Lemma 3.2, here we have $\delta(\lambda)=t_{\rho}$, and $Y$ uniformly distributed over $[0,1]$, so that $\alpha=\omega_{d}/(d+1)$, and $\lambda=\rho t_{\rho}$, so that
$\log\lambda=\left(\frac{d}{d+1}\right)\log\rho+\frac{\log\log\rho+\log(d/\omega_{d})}{d+1}+o(1),$
and $\log\log\lambda=\log\log\rho+\log(d/(d+1))+o(1)$. Hence,
$\alpha\delta(\lambda)^{d}\lambda-\log\lambda-(d-1)\log\log\lambda$ $=\frac{\omega_{d}\rho t_{\rho}^{d+1}}{d+1}-\frac{d\log\rho}{d+1}-\frac{d^{2}\log\log\rho}{d+1}-\frac{\log(d/\omega_{d})}{d+1}$
$~{}~{}~{}~{}~{}~{}~{}-(d-1)\log\left(\frac{d}{d+1}\right)+o(1),$
and by (4.1), as $\rho\to\infty$ this tends to
$\frac{\beta}{d+1}-\log\left(\left(\frac{d^{d^{2}}}{\omega_{d}(d+1)^{d^{2}-1}}\right)^{1/(d+1)}\right).$
Hence we have the condition (3.2) from Lemma 3.2 (with $\beta$ on the right hand side replaced by the last displayed expression). Also $\mathbb{E}\,[Y^{d-1}]=1/d$ and $\mathbb{E}\,[Y^{d}]=1/(d+1)$.
We have $\tilde{T}_{\rho}\leq t_{\rho}$ if and only if $A\subset\cup_{i\in\mathbb{N}}B(X_{i},t_{\rho}-S_{i})$. Therefore by taking $k=1$ and $B_{\lambda}=A$ for all $\lambda$ in Lemma 3.2, we obtain that
$\mathbb{P}[\tilde{T}_{\rho}\leq t_{\rho}]\to\exp\left(-c_{d}\left(\frac{(d+1)^{d-1}}{d^{d}}\right)\left(\frac{d^{d^{2}}}{\omega_{d}(d+1)^{d^{2}-1}}\right)^{1/(d+1)}|A|e^{-\beta/(d+1)}\right).$
This yields (2.11).
For (2.12), take $\rho=L^{d+1}$, so that $\log\log L=\log\log\rho-\log(d+1)$. By Lemma 4.1, $L\tilde{T}_{\rho}$ has the same distribution as $\tilde{\tau}_{L}$. Therefore
$\mathbb{P}\left[\omega_{d}\tilde{\tau}_{L}^{d+1}-d(d+1)\log L-d^{2}\log\log L\leq\beta\right]$
$=\mathbb{P}\left[\omega_{d}\rho\tilde{T}_{\rho}^{d+1}-d\log\rho-d^{2}\log\log\rho+d^{2}\log(d+1)\leq\beta\right],$
and by (2.11), this tends to $\exp(-c_{d}(d^{d}\omega_{d})^{-1/(d+1)}(d+1)^{d^{2}/(d+1)}|A|e^{-\beta/(d+1)}),$ and then (2.12) follows. ∎
We now derive Theorems 2.9 and 2.10 from Theorems 2.1 and 2.2, respectively.
Proof of Theorem 2.9.
We take $Y$ to be uniformly distributed on $[0,1]$. If the restricted J-M model with intensity $\rho$ runs for time $t$, the intensity of the resulting restricted SPBM is $n=t\rho$, and the distribution of radii is that of $tY$. Let $\beta\in\mathbb{R}$, and define
$t_{\rho}:=\Big{(}\frac{(2\log\rho+4\log\log\rho+\beta)\vee 0}{\pi\rho}\Big{)}^{1/3},~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\rho>1,$
so that $\pi\rho t_{\rho}^{3}-2\log\rho-4\log\log\rho=\beta,$ for all large enough $\rho$. Then as $\rho\to\infty$,
$\log(\rho t_{\rho})$ $=(2/3)\log\rho+(1/3)\log\log\rho+\log((2/\pi)^{1/3})+o(1),$
and $\log\log(\rho t_{\rho})=\log\log\rho+\log(2/3)+o(1).$ Therefore since $\mathbb{E}\,[Y^{2}]=1/3$, we have
$\pi\rho t_{\rho}^{3}\mathbb{E}\,[Y^{2}]-\log(\rho t_{\rho})-\log\log(\rho t_{\rho})=$ $\beta/3+(2/3)\log\rho+(4/3)\log\log\rho$
$-(2/3)\log\rho-(1/3)\log\log\rho-\log((2/\pi)^{1/3})$
$-\log\log\rho-\log(2/3)+o(1)$
$=$ $(\beta/3)-\log((16/(27\pi))^{1/3})+o(1),$
so applying Theorem 2.1 with $n=\rho t_{\rho}$, $r_{n}=t_{\rho}$, $k=1$ and $Y$ uniform on $[0,1]$ yields
$\mathbb{P}[\pi\rho T_{\rho}^{3}-2\log\rho-4\log\log\rho\leq\beta]=\mathbb{P}[T_{\rho}\leq t_{\rho}]=\mathbb{P}[A\subset Z_{A,1}(n,r_{n}Y)]$
$\to\exp\Big{(}-\Big{(}\frac{3}{4}\Big{)}|A|\Big{(}\frac{16}{27\pi}\Big{)}^{1/3}e^{-\beta/3}-\Big{(}\frac{3}{4\pi}\Big{)}^{1/2}|\partial A|\Big{(}\frac{16}{27\pi}\Big{)}^{1/6}e^{-\beta/6}\Big{)},$
and hence (2.13).
For (2.14), let $\rho=L^{3}$, so that $\log\log\rho=\log\log L+\log(3)$. Then by Lemma 4.1,
$\mathbb{P}[\pi\tau_{L}^{3}-6\log L-4\log\log L\leq\beta]=\mathbb{P}[\pi\rho T_{\rho}^{3}-2\log\rho-4\log\log\rho+\log 81\leq\beta],$
and by (2.13), this converges to $\exp(-(4\pi)^{-1/3}|A|3^{4/3}e^{-\beta/3}-(2\pi^{2})^{-1/3}|\partial A|3^{4/6}e^{-\beta/6}),$ and hence (2.14). ∎
Proof of Theorem 2.10.
Let $\beta\in\mathbb{R}$ and for $\rho>1$ define
$t_{\rho}:=\Big{(}\frac{(2(d-1)\log\rho+2d(d-1)\log\log\rho+\beta)\vee 0}{\omega_{d}\rho}\Big{)}^{1/(d+1)},$ (4.2)
so that $\omega_{d}\rho t_{\rho}^{d+1}-2(d-1)\log\rho-2d(d-1)\log\log\rho=\beta,$ for all large $\rho$. Then as $\rho\to\infty$,
$\log(\rho t_{\rho})=$ $\frac{d\log\rho+\log\log\rho}{d+1}+\log\Big{(}\Big{(}\frac{2(d-1)}{\omega_{d}}\Big{)}^{1/(d+1)}\Big{)}+o(1),$
and $\log\log(\rho t_{\rho})=\log\log\rho+\log(d/(d+1))+o(1).$ Therefore taking $Y$ to be uniformly distributed on $[0,1]$, so that $\mathbb{E}\,[Y^{d}]=1/(d+1)$, using (4.2) we obtain that
$\rho t_{\rho}\omega_{d}t_{\rho}^{d}\mathbb{E}\,[Y^{d}]-(2-2/d)\log(\rho t_{\rho})-2(d-2+1/d)\log\log(\rho t_{\rho})$
$=\frac{2(d-1)\log\rho+2d(d-1)\log\log\rho}{d+1}+\frac{\beta}{d+1}$
$-\frac{(2-2/d)(d\log\rho+\log\log\rho)}{d+1}-(2-2/d)\log\Big{(}\Big{(}\frac{2(d-1)}{\omega_{d}}\Big{)}^{1/(d+1)}\Big{)}$
$-2(d-2+1/d)\log\log\rho-2(d-2+1/d)\log(d/(d+1))+o(1)$
$=\frac{2d(d-1)-(2-2/d)-2(d-2+1/d)(d+1)}{d+1}\log\log\rho+\frac{\beta}{d+1}-c_{d}^{\prime\prime\prime}+o(1)$
$=\frac{\beta}{d+1}-c_{d}^{\prime\prime\prime}+o(1),$
where we set
$c^{\prime\prime\prime}_{d}:=\log\Big{(}\Big{(}\frac{2(d-1)}{\omega_{d}}\Big{)}^{(2-2/d)/(d+1)}\Big{(}\frac{d}{d+1}\Big{)}^{2(d-2+1/d)}\Big{)}.$
Therefore applying Theorem 2.2 with $k=1$, $n=\rho t_{\rho}$ and $r_{n}=t_{\rho}$, we obtain that if $d\geq 3$ then
$\mathbb{P}[\omega_{d}\rho T_{\rho}^{d+1}-2(d-1)\log\rho-2d(d-1)\log\log\rho\leq\beta]=\mathbb{P}[T_{\rho}\leq t_{\rho}]=\mathbb{P}[A\subset Z_{A,1}(n,r_{n}Y)]$
$\to\exp\Big{(}-\Big{(}\frac{c_{d,1}(d+1)^{d-2+1/d}|\partial A|}{d^{d-1}}\Big{)}\Big{(}\frac{2(d-1)}{\omega_{d}}\Big{)}^{(1-1/d)/(d+1)}$
×$\times\Big{(}\frac{d}{d+1}\Big{)}^{d-2+1/d}e^{-\beta/(2d+2)}\Big{)}$
$=\exp\Big{(}-\frac{c_{d-1}\omega_{d-1}^{2d-3}(d-1)^{d-2+(1/d)+((1-1/d)/(d+1))}|\partial A|e^{-\beta/(2d+2)}}{\omega_{d-2}^{d-1}\omega_{d}^{d-2+(1/d)+((d-1)/(d(d+1)))}2^{1-(1/d)-(1-1/d)/(d+1)}d^{d-1}}\Big{)}$
$=\exp\Big{(}-\frac{c_{d-1}\omega_{d-1}^{2d-3}(d-1)^{d(d-1)/(d+1)}|\partial A|e^{-\beta/(2d+2)}}{\omega_{d-2}^{d-1}\omega_{d}^{d(d-1)/(d+1)}2^{(d-1)/(d+1)}d^{d-1}}\Big{)},$
and hence (2.16).
If $d=2$ then our $t_{\rho}$ defined at (4.2) is the same as in the proof of Theorem 2.9. Applying the case $d=2$ of Theorem 2.2, in the same way as we applied Theorem 2.1 in the proof of Theorem 2.9, gives us the same outcome as in Theorem 2.9, namely (2.13).
When $d=2$ we obtain (2.14) from (2.13) as in the proof of Theorem 2.9. When $d\geq 3$, to get (2.17) we set $\rho=L^{d+1}$. By Lemma 4.1 and (2.16),
$\mathbb{P}[\omega_{d}\tau_{L}^{d+1}-2(d^{2}-1)\log L-2d(d-1)\log\log L\leq\beta]$
$=\mathbb{P}[\omega_{d}\rho T_{\rho}^{d+1}-2(d-1)\log\rho-2d(d-1)[\log\log\rho-\log(d+1)]\leq\beta]$
$\to\exp(-c^{\prime}_{d}|\partial A|e^{-\beta/(2d+2)}(d+1)^{d(d-1)/(d+1)}),$
i.e. (2.17). ∎
Appendix A Confirming consistency with Chiu’s result
In this appendix we verify that our Proposition 2.8 is consistent with [1, Theorem 4]. The latter result is concerned with a quantity denoted $T_{L}$ in [1] which is the same as our $\tilde{\tau}_{L}$, assuming we take $v=\gamma=1$ in [1]. Chiu takes $A=[0,1]^{d}$ and defines $c=c(L)=d(d+1)\log L-\log\omega_{d}$. Then in our notation, the result in [1] says that for any $u\in\mathbb{R}$, as $L\to\infty$ we have
$\mathbb{P}\left[c^{d/(d+1)}\omega_{d}^{1/(d+1)}\tilde{\tau}_{L}-c-\log\left(c^{1/(d+1)}\left(\frac{c+\log c}{d+1}\right)^{d-1}\right)\leq u\right]\to F(u),$ (A.1)
with $F(u)=\exp(-c_{d}(d+1)^{d-1}d^{-d}e^{-u})$ (recall that our $c_{d}$ is the same as Chiu’s $\psi_{d}$ so $c$ and $c_{d}$ are two different things here).
To verify that we can recover (A.1) from Proposition 2.8, note first that $c\to\infty$ as $L\to\infty$ and $\log c=\log\log L+\log(d(d+1))+o(1)$. Also
$\log\left(c^{1/(d+1)}\left(\frac{c+\log c}{d+1}\right)^{d-1}\right)=\log\left(c^{d^{2}/(d+1)}(d+1)^{1-d}\right)+o(1).$
Hence for some $z_{L}$ which tends to zero as $L\to\infty$, the left hand side of (A.1) equals
$\mathbb{P}\left[c^{d}\omega_{d}\tilde{\tau}_{L}^{d+1}\leq\left(c+\log(c^{d^{2}/(d+1)}(d+1)^{1-d})+u+z_{L}\right)^{d+1}\right].$
By binomial expansion, for some $y_{L}$ which tends to zero as $L\to\infty$, this probability equals
$\mathbb{P}\left[c^{d}\omega_{d}\tilde{\tau}_{L}^{d+1}\leq c^{d+1}+(d+1)c^{d}\left(\log(c^{d^{2}/(d+1)}(d+1)^{1-d})+u+z_{L}\right)+y_{L}c^{d}\right]$
$=$ $\mathbb{P}\left[\omega_{d}\tilde{\tau}_{L}^{d+1}\leq c+(d+1)\left(\log(c^{d^{2}/(d+1)}(d+1)^{1-d})+u+z_{L}\right)+y_{L}\right]$
$=$ $\mathbb{P}\left[\omega_{d}\tilde{\tau}_{L}^{d+1}\leq d(d+1)\log L-\log\omega_{d}+d^{2}\log c+\log((d+1)^{1-d^{2}})+(d+1)u+z^{\prime}_{L}\right],$
where we set $z^{\prime}_{L}:=(d+1)z_{L}+y_{L}=o(1)$ as $L\to\infty$. For some $z^{\prime\prime}_{L}$ tending to zero, this probability equals
$\mathbb{P}\left[\omega_{d}\tilde{\tau}_{L}^{d+1}\leq d(d+1)\log L+d^{2}\log\log L+\log\left(\frac{d^{d^{2}}(d+1)^{d^{2}}}{(d+1)^{d^{2}-1}\omega_{d}}\right)+(d+1)u+z^{\prime\prime}_{L}\right],$
$=\mathbb{P}\left[\omega_{d}\tilde{\tau}_{L}^{d+1}-d(d+1)\log L-d^{2}\log\log L\leq\log\left(d^{d^{2}}(d+1)/\omega_{d}\right)+(d+1)u+z^{\prime\prime}_{L}\right].$
By (2.12) from Proposition 2.8, and the continuity of the limiting cumulative distribution function in that result, this probability tends to
$\exp\left(-c_{d}\left(\frac{(d+1)^{d^{2}}}{d^{d}\omega_{d}}\right)^{1/(d+1)}e^{-u}(d^{d^{2}}(d+1))^{-1/(d+1)}\omega_{d}^{1/(d+1)}\right)$
$=\exp\left(-c_{d}d^{-d}(d+1)^{d-1}e^{-u}\right)=F(u).$
In other words, we have derived (A.1) from Proposition 2.8.
Acknowledgment. We thank Edward Crane for a useful conversation on the link between the the SPBM and the Johnson-Mehl model.